### v1

Abstract
A $(\phi,\epsilon)$-expander-decomposition of a graph $G$ (with $n$ vertices and $m$ edges) is a partition of $V$ into clusters $V_{1},\ldots,V_{k}$ with conductance $\Phi(G[V_{i}])\geq\phi$, such that there are at most $\epsilon m$ inter-cluster edges. Such a decomposition plays a crucial role in many graph algorithms. [2] gave a randomized $\tilde{O}(m)$ time algorithm for computing a $(\phi,\phi\log^{2}{n})$-expander decomposition.
In this paper we generalize this result for a broader notion of expansion. Let $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$ be a vertex measure. A standard generalization of conductance of a cut $(S,\bar{S})$ is its $\mu$-expansion $\Phi^{\mu}_{G}(S,\bar{S})=|E(S,\bar{S})|/\min\{\mu(S)),\mu(\bar{S})\}$, where $\mu(S)=\sum_{v\in S}\mu(v)$.
We present a randomized $\tilde{O}(m)$ time algorithm for computing a $(\phi,\phi\log^{2}{n}\left(\frac{\mu(V)}{m}\right))$-expander decomposition with respect to $\mu$-expansion.
1 Introduction
The conductance of a cut $(S,\bar{S})$ is $\Phi_{G}(S,\bar{S})=\frac{|E(S,\bar{S})|}{\min(\operatorname{\mathbf{vol}}(S),\operatorname{\mathbf{vol}}(\bar{S}))}$, where $\operatorname{\mathbf{vol}}(S)$ is the sum of the degrees of the vertices of $S$. The conductance of a graph $G$ is the smallest conductance of a cut in $G$.
A $(\phi,\epsilon)$-expander decomposition of a graph $G$ is a partition of the vertices of $G$ into clusters $V_{1},\ldots,V_{k}$, each with conductance $\Phi(G[V_{i}])\geq\phi$ such that there are at most $\epsilon m$ inter-cluster edges, where $\phi,\epsilon\geq 0$. Consider the problem of computing a $(\phi,\epsilon)$-expander decomposition for a given graph $G$ and $\phi>0$, while minimizing $\epsilon$ as a function of $\phi$. It is known that a $(\phi,\epsilon)$-expander decomposition, with $\epsilon=O(\phi\log n)$, always exists and that $\epsilon=\Theta(\phi\log n)$ is best possible [12, 1]. In a recent result, [2] achieved a $(\phi,\phi\log^{2}{n})$-expander decomposition in $\tilde{O}(m)$ time.
The following is an overview on the result of [2] which builds on [12]. The naive way to compute an expander decomposition is to use an approximation algorithm for the sparsest cut problem. Given an $f(n)$-approximation algorithm for the problem of finding a minimum conductance cut, one can get a $(\phi,O(f(n)\cdot\phi\log n))$-expander decomposition algorithm by recursively computing approximate cuts until all components are certified as expanders. The drawback of this approach is a possible linear recursion depth (because of unbalanced cuts - one side has very small volume) which affects the running time. To solve this problem, Saranurak and Wang [12] developed a sparse cut algorithm that either returns a balanced sparse cut or an unbalanced sparse cut in which the larger side is an expander (and therefore they only need to recur on the smaller side of the cut). Their algorithm is based on the Cut-Matching Game [5].
Cut-matching: The cut-matching game, played by a cut player and a matching player to construct an expander, is a well known technique to reduce the approximation task for sparsest cut to a polylogarithmic number of maximum flow problems [5]. Khandekar et al. [5] devised a strategy for the cut player that translates to an $O(\log^{2}{n})$ approximation algorithm for the sparsest cut problem. This cut player of [5] is robust in the sense that it is simple and flexible to modifications for other applications [10, 12]. A more sophisticated spectral cut player was devised by [9] which yields $O(\log{n})$ approximation. This cut player is less flexible and harder to modify.
By modifying the cut-matching game and the cut player of [5], Saranrurak and Wang [12] achieved a $(\phi,\phi\log^{3}{n})$-expander decomposition in $\tilde{O}(m)$ time. In a recent result [2] were able by a careful spectral analysis, to combine the techniques of [12] and the spectral cut player of [9] to get a $(\phi,\phi\log^{2}{n})$-expander decomposition in $\tilde{O}(m)$ time.
A useful generalization of conductance is $\mu$-expansion. Let $\mu\in V\rightarrow\mathop{\mathbb{R}}_{\geq 0}^{n}$ be a vertex measure. The $\mu$-expansion of a cut $(S,\bar{S})$ is $\Phi^{\mu}_{G}(S,\bar{S})=\frac{|E(S,\bar{S})|}{\min(\mu(S),\mu(\bar{S}))}$, where $\mu(S)=\sum_{v\in S}\mu(v)$. If $\min(\mu(S),\mu(\bar{S}))=0$ then $\Phi^{\mu}_{G}(S,\bar{S})=\infty$. Note that by setting $\mu(v)=deg_{G}(v)$, we get conductance. The $\mu$-expansion of $G$ is defined as $\Phi^{\mu}(G)=\min_{S\subseteq V}\Phi^{\mu}_{G}(S,\bar{S})$. We say that $G$ is a $(\phi,\mu)$-expander if $\Phi^{\mu}(G)\geq\phi$.
A $(\phi,\epsilon,\mu)$-expander decomposition of a graph $G$ is a partition of the vertices of $G$ into clusters $V_{1},\ldots,V_{k}$, each with $\mu$-expansion $\Phi^{\mu}(G[V_{i}])\geq\phi$ such that there are at most $\epsilon m$ inter-cluster edges, where $\phi,\epsilon\geq 0$ and $\mu\in V\rightarrow\mathop{\mathbb{R}}_{\geq 0}^{n}$.
Our Contribution: In this paper we generalize the result of [2] to the notion of $\mu$-expansion and present an $\tilde{O}(m)$ algorithm for computing a $(\phi,\phi\log^{2}{n}\frac{\mu(V)}{m})$-expander decomposition with respect to $\mu$-expansion. The generalization is quite straightforward; we derive it by modifying a simplified version of  [2] (This simplification is possible due to a recent result on fair-cut [7]). As an immediate consequence of our balanced sparse cut procedure, used in the expander decomposition algorithm (see Lemma 3.1), we get an $O(\log{n})$ approximation algorithm to the sparsest cut problem with respect to a vertex measure $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$.
Related work: In a recent paper, Ameranis et al. [3] use a generalized notion of $\mu$-expansion, also mentioned in [8], where they allowed to consider overlapping partitions of $V$. They define a corresponding generalized version of the cut-matching game, and show how to use a cut strategy for this game to get an approximation algorithm for two generalized cut problems. They claim that one can construct a cut strategy for this measure using ideas from [8]. [^1] The details of such a cut player do not appear in [3] or [8].
The structure of this paper is as follows. Section 2 contains basic definitions. In section 3 we present a version of the cut-matching game, together with a cut player and a matching player which are suitable for our expander decomposition algorithm with respect to a vertex measure $\mu$. Finally, Section 4.3 contains the balanced cut algorithm, from which we obtain an expander decomposition in a standard way which is described in Appendix 5. Appendix A contains missing proofs. Appendices B and C contain algebraic tools and probabilistic lemmas, respectively.
2 Preliminaries
Throughout the paper we work with the undirected graph $G=(V,E)$. We denote the transpose of a vector or a matrix $x$ by $x^{\prime}$. That is if $v$ is a column vector then $v^{\prime}$ is the corresponding row vector. For a vector $v\in\mathop{\mathbb{R}}^{n}$, define $\sqrt{v}$ to be vector whose coordinates are the square roots of those of $v$. Given $A\in\mathop{\mathbb{R}}^{n\times n}$, we denote by $A(i,j)$ the element at the $i$’th row and $j$’th column of $A$. We denote by $A(i,),A(,i)$ the $i$’th row and column of $A$, respectively. We think of both $A(i,)$ and $A(,i)$ as column vectors. We use the abbreviation $A(i)\mathrel{\mathop{:}}=A(i,)$ only with respect to the rows of $A$. Given a vector $v\in\mathop{\mathbb{R}}^{n}$, we denote its $i$’th element by $v(i)$. For disjoint $A,B\subseteq V$, we denote by $E_{G}(A,B)$ the set of edges connecting $A$ and $B$. We denote by $|E_{G}(A,B)|$ the number of edges in $E_{G}(A,B)$, or the sum of their weights if the graph is weighted. We sometimes omit the subscript when the graph is clear from the context. For a weighted graph $H=(V,E,w)$, we define $deg_{H}(v)=w\left(E_{H}(\{v\},V\setminus\{v\})\right)$. If $A=V\setminus B$, then we call $(A,B)$ a cut. Throughout the paper we work with the vertex measure $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$. For $S\subseteq V$ we let $\mu(S)=\sum_{v\in S}\mu(v)$. We define the set of terminals to be $\mathcal{T}=\{v\in V\mid\mu(v)>0\}$. We assume the following on $\mu$.
Assumption 2.1.
For every $v\in\mathcal{T}$ we have $1/poly(n)\leq\mu(v)\leq poly(n)$.
Definition 2.2 ($\mu$-Matching).
Given a vector $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$ and a collection of weighted pairs $M=\left\{(u_{i},v_{i},w_{i})\right\}_{i=1}^{m}$. We say that $M$ is a $\mu$-matching if the graph defined by $M$ satisfies $deg_{M}(v)=\mu(v)$, for every $v$.
Definition 2.3 ($\mu$-stochastic).
A matrix $F\in\mathop{\mathbb{R}}^{n\times n}$ is $\mu$-stochastic with respect to $\mu\in\mathop{\mathbb{R}}^{n}$ if the following two conditions hold: (1) $F\cdot\mathbbm{1}_{n}=\mu$ and (2) $\mathbbm{1}_{n}^{\prime}\cdot F=\mu^{\prime}$.
Definition 2.4 (Laplacian, Normalized Laplacian).
Let $A\in\mathop{\mathbb{R}}^{n\times n}$ be a symmetric matrix and let $d=A\cdot\mathbbm{1}_{n},\;D=\operatorname{\mathbf{diag}}(d)$. The Laplacian of $A$ is defined as $\mathcal{L}(A)=D-A$. The normalized-Laplacian of $A$ is defined as $\mathcal{N}(A)=D^{-\frac{1}{2}}\mathcal{L}(A)D^{-\frac{1}{2}}=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$. The (normalized) Laplacian of an undirected graph is defined analogously using its adjacency matrix.
Definition 2.5 ($\mu$-expansion).
Let $G=(V,E)$ and $\subset V, SS\neq\emptyset$. Let $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$. The $\mu$-expansion of the cut $(S,\bar{S})$, denoted by $\Phi^{\mu}_{G}(S,\bar{S})$, is
$\Phi^{\mu}_{G}(S,\bar{S})=\frac{|E(S,\bar{S})|}{\min(\mu(S),\mu(\bar{S}))}.$
The $\mu$-expansion of $G$ is defined to be $\Phi^{\mu}(G)=\min_{S\subseteq V}\Phi_{G}(S,\bar{S})$.
Remark 2.6.
In the above definition we allow the denominator to be 0, in which case the $\mu$-expansion of the cut is $\infty$. Therefore, the minimization in the definition of $\Phi^{\mu}(G)$ can be taken over sets $S\subseteq V$ satisfying $\mu(S),\mu(\bar{S})>0$.
Definition 2.7 ($\mu$-Expander, $\mu$-Near-Expander).
Let $G=(V,E)$. We say that $G$ is a $(\phi,\mu)$-expander if $\Phi^{\mu}(G)\geq\phi$. Let $A\subseteq V$. We say that $A$ is a near $(\phi,\mu)$-expander in $G$ if
$\min_{S\subseteq A}\frac{|E(S,V\setminus S)|}{\min(\mu(S),\mu(A\setminus S))}\geq\phi.$
That is, a near expander is allowed to use cut edges that go outside of $A$. Note that a subset of a near expander is also a near expander. This notion is useful in several contexts. Specifically, using “trimming” of [12], under certain conditions on $A$, one can extract an expander $B\subseteq A$ by “trimming” a part of $A$ (see Theorem 5.1).
Definition 2.8 (Embedding).
Let $G=(V,E)$ be an undirected graph. Let $F\in\mathop{\mathbb{R}}^{V\times V}_{\geq 0}$ be a matrix (not necessarily symmetric). We say that $F$ is embeddable in $G$ with congestion $c$, if there exists a multi-commodity flow $f$ in $G$, with $|V|$ commodities, one for each vertex (the vertex $v$ is the source of its commodity), such that, simultaneously for each $(u,v)\in V\times V$, $f$ routes $F(u,v)$ units of $u$’s commodity from $u$ to $v$, and the total flow on each edge is at most $c$.
If $F$ is the weighted adjacency matrix of a graph $H$ on the same vertex set $V$, we say that $H$ is embeddable in $G$ with congestion $c$ if $F$ is embeddable in $G$ with congestion cc. [^2] This definition requires to route $F(u,v)=F(v,u)$ both from $u$ to $v$ and from $v$ to $u$ if $F$ is symmetric.
Lemma 2.9.
Let $G,H$ be two graphs on the same vertex set $V$. Let $A\subseteq V$. Assume that $H$ is embeddable in $G$ with congestion $c$, and that $A$ is a near $(\phi,\mu)$-expander in $H$. Then, $A$ is a near $(\frac{2\phi}{c},\mu)$-expander in $G$.
Proof.
Let $S\subseteq A,\bar{S}=V\setminus S$ be a cut, and assume $\mu(S)\leq\mu(A\setminus S)$. In the embedding of $H$ in $G$, each edge $(u,v)\in E(H)$ corresponds to $w_{H}(u,v)$ units of flow, routed in $G$ from $u$ to $v$. Since each edge of $G$ carries at most $c$ units of flow we get that $|E_{G}(S,\bar{S})|\geq\frac{2}{c}|E_{H}(S,\bar{S})|$. [^3] Recall that by definition for every $\{u,v\}\in E(H)$ we send flow in $G$ in both directions. Since $A$ is a near $(\phi,\mu)$-expander in $H$, we get that
$\Phi^{\mu}_{G}(S,\bar{S})=\frac{|E_{G}(S,\bar{S})|}{\mu(S)}\geq\frac{2}{c}\frac{|E_{H}(S,\bar{S})|}{\mu(S)}\geq\frac{2\phi}{c}.$
Corollary 2.10.
Let $G,H$ be two graphs on the same vertex set $V$. Assume that $H$ is embeddable in $G$ with congestion $c$, and that $H$ is a $(\phi,\mu)$-expander. Then, $G$ is a $(\frac{2\phi}{c},\mu)$-expander.
Proof.
This follows from Lemma 2.9 by choosing $A=V$. ∎
3 Balanced Sparse-Cut via spectral Cut-Matching
In this section we state our main theorem and present our algorithm, whose analysis appears in the next section. The result of [2] for finding a balance sparse cut for the case of conductance is a special case of our theorem in which the vertex measure corresponds to degrees (i.e. $\mu(v)=deg_{G}(v)$ for every $v\in V$).
Theorem 3.1.
Given a graph $G=(V,E)$, a vertex measure $\mu\in\mathop{\mathbb{R}}_{\geq 0}^{n}$ that satisfies Assumption 2.1 and a parameter $\phi>0$, there exists a randomized algorithm which takes $\tilde{O}\left(m\right)$ time [^4] The $\tilde{O}$ hides $\log^{c}n$ factors as well as $\log^{c}\mu^{\star}$ factor, where $\mu^{\star}=\frac{\max(\mu(\cdot))}{\min(\mu(\cdot))}$. and must end in one of the following three cases:
  1. 1.
    We certify that $G$ has $\mu$-expansion $\Phi^{\mu}(G)=\Omega(\phi)$ with high probability.
  2. 2.
    We find a cut $(R,A)$ in $G$ of $\mu$-expansion $\Phi^{\mu}_{G}(R,A)=O(\phi\log n)$, and $\mu(R),\mu(A)$ are both $\Omega(\frac{\mu(V)}{\log n})$, i.e, we find a relatively balanced low conductance cut.
  3. 3.
    We find a cut $(R,A)$ with $\Phi^{\mu}_{G}(R,A)\leq c_{0}\phi\log n$ for some constant $c_{0}$, and $\mu(R)\leq\frac{\mu(V)}{10c_{0}\log n}$, and with high probability $A$ is a near $(\Omega(\phi),\mu)$-expander in $G$.
This theorem is a straightforward generalization of Theorem $5.1$ of [2]. The main algorithmic difference is the usage of fractional matchings during the cut matching game and zeroing out in the random walk matrix rows/columns corresponding to vertices $v\notin\mathcal{T}$. (Recall that $\mathcal{T}=\{v\in V\mid\mu(v)>0\}$.)
To get Theorem 3.1 we use the cut matching game, specified in the next subsection. Theorem 3.1 implies the following theorem.
Theorem 3.2.
Given a graph $G=(V,E)$ of $m$ edges and a parameter $\phi$, there is a randomized algorithm that with high probability finds a partition of $V$ into clusters $V_{1},...,V_{k}$ such that $\forall i:\Phi^{\mu}_{G[V_{i}]}=\Omega(\phi)$ and $\sum_{i}{|E(V_{i},\bar{V_{i}})|}=O(\phi\mu(V)\log^{2}n)$. [^5] If we remove Assumption 2.1, the exact bound is $\sum_{i}{|E(V_{i},\bar{V_{i}})|}=O(\phi\mu(V)\log n\log\mu^{\star})$, where $\mu^{\star}=\frac{\max(\mu(\cdot))}{\min(\mu(\cdot))}$. The running time of the algorithm is $\tilde{O}(m)$.
The derivation of Theorem 3.2 using Theorem 3.1 uses the same methods as in [12, 7]. For completeness, we give this derivation in Appendix 5.
The rest of this section is organized as follows. In Subsection 3.1 we introduce a generalized version of the cut matching game of [2]. In Subsections 3.2 and 3.3 we present a cut player and a matching player, respectively. Finally, in Subsection 3.4 we present the algorithm for Theorem 3.2 using the cut and matching players.
3.1 Cut matching for $\mu$-expansion
The following is a generalized version of the cut matching games in [12, 2], suited for $\mu$.
The Cut-Matching game for $\mu$-expansion, with parameters $T$ and a vector $\mu$: The game is played on a series of graphs $G_{i}$. Initially, $G_{0}=\emptyset,A_{0}=V$. At iteration $t$, the cut player produces two weighted sets, $L_{t},R_{t}\subseteq A_{t}$, with weights $m_{v},\bar{m}_{v}$, respectively. The weights satisfy $m(L_{t})\leq\mu(A_{t})/8,\bar{m}(R_{t})\geq\mu(A_{t})/2$, and the total weight $m_{v}+\bar{m}_{v}$ of each $v\in V$ in $L_{t}\cup R_{t}$ is at most $\mu(v)$. The matching player responds with a set $S_{t}\subseteq A_{t}$ and a weighted matching $M_{t}$ that only matches vertices in $L_{t}\setminus S_{t}$ to vertices in $R_{t}\setminus S_{t}$. We set $G_{t+1}=G_{t}\cup M_{t},A_{t+1}=A_{t}\setminus S_{t}$. The union $G_{t}\cup M_{t}$ refers to the graph whose weighted adjacency matrix is the sum of the weighted adjacency matrices of $G_{t}$ and $M_{t}$. The game ends at iteration $T$, and the quality of the game is $r$, the near $\mu$-expansion of $A_{t}$ in $G_{t}$.
Our cut matching game iteratively shrinks the domain $A_{t}\subseteq V$. So we start by introducing the following matrices that would be useful to work with a shrinking domain.
Definition 3.3 ($I_{t},\mu_{t},U_{t},P_{t}$).
We define the following variables
  1. 1.
    $I_{t}=diag(\mathbbm{1}_{A_{t}\cap\mathcal{T}})\in\mathop{\mathbb{R}}^{n\times n}$. Similarly, $\mathbbm{1}_{t}=I_{t}\cdot\mathbbm{1}$.
  2. 2.
    $\mu_{t}=I_{t}\cdot\mu\in\mathop{\mathbb{R}}^{n}$, i.e the projection of $\mu$ onto $A_{t}$.
  3. 3.
    $U_{t}=I_{t}\cdot U=\operatorname{\mathbf{diag}}(\mu_{t})\in\mathop{\mathbb{R}}^{n\times n}$, where $U=\operatorname{\mathbf{diag}}(\mu)$.
  4. 4.
    $P_{t}=I_{t}-\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}\in\mathop{\mathbb{R}}^{n\times n}$. This matrix projects on the orthogonal complement of $\sqrt{\mu_{t}}$.
Remark 3.4.
Throughout the paper we will abuse the inverse notation and write $U^{-1}$ (or $U^{-\frac{1}{2}}$) to the denote the pseudo inverse of $U$ (i.e. inverse each entry on the diagonal if it is not zero).
3.2 Cut player
Let $\delta=\Theta(\log{n})$ be a power of $2$ which we formally set later in the proof of Lemma 4.12.666We need $\delta$ to be a power of $2$ to apply Theorem  B.3 to a $\delta$ matrix power later on. The cut player implicitly maintains a $\mu$-stochastic flow matrix $F_{t}\in\mathop{\mathbb{R}}^{n\times n}$, and the graph $G_{t}$ which is the union of the matchings that it obtained so far from the matching player ($t$ is the index of the round). The flow $F_{t}$ and the graph $G_{t}$ have two crucial properties. First, we can embed $F_{t}$ in $G_{t}$ with $O(\frac{1}{\delta})$ congestion (Lemma 4.4) and embed $G_{t}$ in $G$ with $O(\frac{t}{\phi\log{n}})$ congestion (Lemma 4.5), see Definition 2.8. Second, after $T=\Theta(\log^{2}n)$ rounds, $A_{t}$ will be a near $(\Omega(1),\mu)$-expander in $F_{T}$ with high probability (Lemma 4.12). By Lemma 2.9, this implies that, with hight probability, $A_{t}$ will be a near $(\Omega(\phi),\mu)$-expander in $G$.
We define the matrix $W_{t}=(P_{t}U^{-\frac{1}{2}}F_{t}U^{-\frac{1}{2}}P_{t})^{\delta}$ (we recall again that $\delta=\Theta(\log{n})$ as specified above). This definition makes us “focus” only on the remaining vertices $A_{t}\cap\mathcal{T}$, as any row/column of $W_{t}$ corresponding to a vertex $v\in V\setminus(A_{t}\cap\mathcal{T})$ is zero. The matrix $W_{t}$ is used in this section to define the projections that the cut player needs to update $F_{t}$. It is also used in Section 4.2 to define the potential that measures how far is the remaining part of the graph $A_{T}$ from a near $\mu$-expander. In particular, we show in Lemma 4.12 and Corollary 4.13 that if $W_{T}^{2}$ has small eigenvalues (which will be the case when the potential is small) then $A_{T}$ is near $(\Omega(\delta),\mu)$-expander in $G_{T}$.
The cut player updates $F_{t}$ as follows. Let $r\in\mathop{\mathbb{R}}^{n}$ be a random unit vector. Consider the projections $u_{i}=\frac{1}{\sqrt{\mu(i)}}\langle W_{t}(i),r\rangle$, for $i\in A_{t}\cap\mathcal{T}$ and $u_{i}=0$ otherwise. Note that since $P_{t}\sqrt{\mu_{t}}=0$, and $W_{t}$ is symmetric we have that:
$\sum_{i\in A_{t}}{\mu(i)u_{i}}=\sum_{i\in A_{t}}{\sqrt{\mu(i)}\left\langle W_{t}(i),r\right\rangle}=\left\langle\sum_{i\in A_{t}}{\sqrt{\mu(i)}W_{t}(i)},r\right\rangle=\left\langle\left(\sqrt{\mu^{\prime}_{t}}W_{t}\right)^{\prime},r\right\rangle=\left\langle W_{t}\sqrt{\mu},r\right\rangle=0$
We use the following lemma to partition (some of) the vertices of $A_{t}$ into two weighted sets $A_{t}^{l}$ and $A_{t}^{r}$. [^7] Note that this does not produce a bisection of $A_{t}$. The lemma is a continuous modification of Lemma 3.3 in [10]. For completeness, we include the proof in Appendix A.
Lemma 3.5 (Lemma 3.3 in [10]).
Given $u_{i}\in\mathop{\mathbb{R}}$ for all $i\in A_{t}$, such that $\sum_{i\in A_{t}}{\mu(i)u_{i}}=0$, we can find in time $O(|A_{t}|\cdot\log(|A_{t}|))$ a set of source nodes $A^{l}_{t}\subseteq A_{t}$, a set of target nodes $A^{r}_{t}\subseteq A_{t}$, weights $0\leq m_{i},\bar{m}_{i}\leq\mu(i)$ for $i\in A_{t}$, and a separation value $\eta$ such that:
  1. 1.
    $\eta$ separates the sets $A^{l}_{t},A^{r}_{t}$, i.e., either $\max_{i\in A^{l}_{t}}{u_{i}}\leq\eta\leq\min_{j\in A^{r}_{t}}{u_{j}}$, or $\min_{i\in A^{l}_{t}}{u_{i}}\geq\eta\geq\max_{j\in A^{r}_{t}}{u_{j}}$,
  2. 2.
    $\forall i\in V:m_{i}+\bar{m}_{i}\leq\mu(i)$,
  3. 3.
    $\bar{m}(A^{r}_{t})\geq\frac{\mu(A_{t})}{2}$, $m(A^{l}_{t})\leq\frac{\mu(A_{t})}{8}$,
  4. 4.
    $\forall i\in A^{l}_{t}:(u_{i}-\eta)^{2}\geq\frac{1}{9}u_{i}^{2}$,
  5. 5.
    $\sum_{i\in A^{l}_{t}}{m_{i}u_{i}^{2}}\geq\frac{1}{80}\sum_{i\in A_{t}}{\mu(i)u_{i}^{2}}$.
Note that a vertex could appear both in $A_{t}^{l}$ and in $A_{t}^{r}$, if $u_{i_{j}}=\eta$. The cut player sends $A_{l},A_{r}$ and $A_{t}$ to the matching player.
In turn, the matching player (see Subsection 3.3) returns a cut $(S_{t},A_{t}\setminus S_{t}$) and a fractional matching $M_{t}$ of total weight $m(A^{l}_{t}\setminus S_{t})$ between $A^{l}_{t}\setminus S_{t}$ and $A^{r}_{t}\setminus S_{t}$. Define $N_{t}=\frac{\delta-1}{\delta}U+\frac{1}{\delta}M_{t}$. The cut player then updates $F_{{t}}$ as follows: $F_{{t+1}}=N_{t}\cdot U^{-1}F_{t}U^{-1}N_{t}$. Intuitively, this applies lazy (with respect to $\delta$) averaging of the rows and columns of $F_{t}$ according to $M_{t}$. For the analysis, we update $G_{t+1}$ as $G_{t+1}=G_{t}\cup M_{t}$. [^8] We allow for parallel edges. Note that $G_{t+1}$ might include self-loops because of $A^{l}_{t}\cap A^{r}_{t}$.
3.3 Matching player
The matching player receives $A^{l}_{t}$ and $A^{r}_{t}$ and the current $A_{t}$. For a vertex $v\in V$, denote by $m_{v}$ the weight of $v$ in $A_{t}^{l}$, and by $\bar{m}_{v}$ the weight of $v$ in $A_{t}^{r}$ ($m_{v}+\bar{m}_{v}\leq\mu(v)$). The matching player solves the flow problem on $G[A_{t}]$, specified by the following lemma, that closely resembles Lemma $8.1$ of [7] when setting $V=A_{t}$. For completeness, we include the proof in Appendix A.2.
Lemma 3.6 (Lemma $8.1$, [7]).
Let $G=(V,E)$ be a graph with $n$ vertices and $m$ edges, let $A^{l},A^{r}\subseteq V$, and let $\phi>0$ be a parameter. For a vertex $v\in V$, denote by $m_{v}$ the weight of $v$ in $A^{l}$, and by $\bar{m}_{v}$ the weight of $v$ in $A^{r}$. Assume that $m_{v}+\bar{m}_{v}\leq\mu(v)$ and that $\bar{m}(A^{r})\geq\frac{1}{2}\mu(V),m(A^{l})\leq\frac{1}{8}\mu(V)$. We define the flow problem $\Pi(G)$, as the problem in which a source $s$ is connected to each vertex $v\in A^{l}$ with an edge of capacity $m_{v}$ and each vertex $v\in A^{r}$ is connected to a sink $t$ with an edge of capacity $\bar{m}_{v}$. Every edge of $G$ has the same capacity $c=\Theta\left(\frac{1}{\phi\log n}\right)$. A feasible flow for $\Pi(G)$ is a maximum flow that saturates all the edges outgoing from $s$. Then, in time $\tilde{O}(m)$, we can find either
  1. 1.
    A feasible flow $f$ for $\Pi(G)$; or
  2. 2.
    A cut $S$ where $\Phi_{G}^{\mu}(S,V\setminus S)\leq\frac{7}{c}=O(\phi\log n)$, $\mu(V\setminus S)\geq\frac{1}{3}\mu(V)$ and a feasible flow for the problem $\Pi(G-S)$, where we only consider the sub-graph $G[V\setminus S\cup\{s,t\}]$ (that is, vertices $v\in A^{l}\setminus S$ are sources of $m_{v}$ units, and vertices $v\in A^{r}\setminus S$ are sinks of $\bar{m}_{v}$ units).
Remark 3.7.
It is possible that $A^{l}\subseteq S$, in which case the feasible flow for $\Pi(G-S)$ is trivial (the total source mass is 0).
Let $S_{t}$ be the cut returned by the lemma. If the lemma terminates with the first case, we set $S_{t}=\emptyset$. Using dynamic trees [11], we can decompose the returned flow into a set of (at most $m$) weighted paths, each carrying some flow from a vertex $u\in A_{t}^{l}\setminus S_{t}$ to a vertex $v\in A^{r}_{t}\setminus S_{t}$. If $u\in A^{l}_{t}\cap A^{r}_{t}$ then it is possible that a path starts and ends at $u$. Each $u\in A^{l}_{t}\setminus S_{t}$ is the starting point of exactly $m_{u}\leq\mu(u)$ flow, and each $v\in A^{r}_{t}\setminus S_{t}$ is the endpoint of at most $\bar{m}_{v}\leq\mu(v)$ flow. Define the weighted matching $\tilde{M_{t}}$ as $\tilde{M_{t}}=((u_{i},v_{i},w_{i}))_{i=1}^{|A^{l}_{t}\setminus S_{t}|}$, where $u_{i}$ and $v_{i}$ are the endpoints of path $i$ which carries $w_{i}$ flow. We can view $\tilde{M_{t}}$ as a symmetric $n\times n$ matrix, such that $\tilde{M_{t}}(u,v)=\tilde{M_{t}}(v,u)=\tilde{w}_{uv}$, where $\tilde{w}_{uv}$ is the total flow sent on paths between $u$ and $v$. We turn $\tilde{M_{t}}$ into a $\mu$-stochastic matrix by increasing its diagonal entries by $\mu-\tilde{M_{t}}\mathbbm{1}_{n}$. Formally, we set $M_{t}\mathrel{\mathop{:}}=\tilde{M_{t}}+\operatorname{\mathbf{diag}}(\mu-\tilde{M_{t}}\mathbbm{1}_{n})$. Notice that $\mu-\tilde{M_{t}}\mathbbm{1}_{n}$ has only non-negative entries, so $M_{t}$ also has non-negative entries. The matching player sends back $M_{t},S_{t}$ as the response to the subsets $A^{l}_{t}$ and $A^{r}_{t}$ given by the cut player.
3.4 The Algorithm
Let $T=\Theta(\log^{2}n)$ and $c=\Theta(\frac{1}{\phi\log n})$. The algorithm for Theorem 3.1 is presented in Algorithm 1. The algorithm runs for at most $T$ rounds and stops when $\mu(R_{t})>\frac{\mu(V)\cdot c\cdot\phi}{70}=\Omega(\frac{\mu(V)}{\log n})$. In each round $t$, we implicitly update $F_{t}$ (see Section 3.2). In order to keep the running time near linear, we do the flow computation using the fair cut result of [7]. These routines may also return a cut $S_{t}\subseteq A_{t}$ with $\Phi_{G[A]}(S_{t},A_{t}\setminus S_{t})\leq\frac{7}{c}$, in which case we “move” $S_{t}$ from $A_{t+1}$ to $R_{t+1}$. After $T$ rounds, $F_{T}$ certifies that the remaining part of $A_{T}$ is a near $(\Omega(\phi),\mu)$-expander.
Algorithm 1 Cut-Matching
1:function Cut-Matching($G,\phi$)
2:  $T\leftarrow\Theta(\log^{2}n)$, $\leftarrow\Theta(\frac{1}{\phi\log n}). \triangleright cc$ is an integer.
3:  $t\leftarrow 0$.
4:  $A_{0}\leftarrow V$,  $R_{0}\leftarrow\emptyset$.
5:  Set $F_{0}\leftarrow U\mathrel{\mathop{:}}=\operatorname{\mathbf{diag}}(\mu)$.
6:  while $\mu(R_{t})\leq\frac{\mu(V)\cdot c\cdot\phi}{70}$ and $t<T$ do
7:   Update $F_{t}$ to $F_{t+1}$. $\triangleright$ Sections 3.2 and 3.3 describe this update.
8:   This update returns $S_{t}\subseteq A_{t}$ where $\Phi^{\mu}_{G[A_{t}]}(S_{t},A_{t}\setminus S_{t})\leq\frac{7}{c}$, or $S_{t}=\emptyset$.
9:   $A_{t+1}\leftarrow A_{t}-S_{t}$, $R_{t+1}\leftarrow R_{t}\cup S_{t}$.
10:   $t\leftarrow t+1$.
11:  end while
12:  if $t=T$ then
13:   if $R=\emptyset$ then
14:     Certify that $\Phi^{\mu}(G)=\Omega(\phi)$. $\triangleright$ Case (1) of Theorem 3.1.
15:   else if $\mu_{G}\{R_{t}\}>\frac{\mu(V)\cdot c\cdot\phi}{70}$ then  return $(A_{T},R_{T})$. $\triangleright$ Case (2) of Theorem 3.1.
16:   else return $(A_{T},R_{T})$. $\triangleright$ Case (3) of Theorem 3.1.
17:   end if
18:  else return $(A_{T},R_{T})$. $\triangleright$ Case (2) of Theorem 3.1.
19:  end if
20:end function
4 Proof of Theorem 3.1
In this section we prove that Algorithm 1 satisfies the conditions of Theorem 3.1. We start by giving a high-level overview of the analysis of our algorithm. The potential we use here to measure the progress of the cut player is $\psi(t)=\operatorname{Tr}[W_{t}^{2}]=\sum_{i\in A_{t}}{\left\|W_{t}(i)\right\|_{2}^{2}}$, where $W_{t}$ was defined as $W_{t}=(P_{t}U^{-\frac{1}{2}}F_{t}U^{-\frac{1}{2}}P_{t})^{\delta}$. Recall $\delta=O(\log n)$ is a power of $2$, which will be defined in the proof of Lemma 4.12.
Our key technical Lemmas 4.8 and 4.9 carefully establish the decrease in potential achieved by the cut player, taking into account the matchings and the deleted cuts. We conclude in Corollary 4.11 that after $O(\log^{2}n)$ iterations the potential falls below $1/n$. We then prove, using our variation on Cheeger’s inequality (see Lemma 4.12), that having a small potential means that $A_{t}$ is a near $(\Omega(1),\mu)$-expander in $F_{t}$.
Using the fact that $F_{t}$ is embeddable in $G_{t}$ (Lemma 4.4) and that $G_{t}$ is embeddable in $G$ (Lemma 4.5) we can conclude that if we reach round $T$, then with high probability, $A_{T}$ is a near $(\Omega(\phi),\mu)$-expander in $G$ (Corollary 4.13).
This section is organized as follows. Subsection 4.1 shows that $F_{t}$ is embeddable in $G_{t}$ with congestion $\frac{2}{\delta}$ and that $G_{t}$ is embeddable in $G$ with congestion $c\cdot t$. Subsection 4.2 shows that if we reach round $T$, then with high probability, $A_{T}$ is a near $(\Omega(\phi),\mu)$-expander in $G$. Finally, in Subsection 4.3 we prove Theorem 3.1.
4.1 $F_{t}$ is embeddable in $G$
To begin the analysis of the algorithm, we first define a blocked matrix. This notion will be useful when our matrices “operate” only on vertices of $A_{t}$.
Definition 4.1.
Let $A\subseteq V$. A matrix $B\in\mathop{\mathbb{R}}^{n\times n}$ is $A$-blocked if $B(i,j)=0$ for all $i\neq j$ such that $(i,j)\notin A\times A$. Diagonal entries may be non-zero.
Lemma 4.2.
The following holds for all $t$:
  1. 1.
    $M_{t},N_{t},F_{t}$ and $W_{t}$ are symmetric.
  2. 2.
    $M_{t}$ and $N_{t}$ are $A_{t}\cap\mathcal{T}$-blocked. $F_{t}$ is $\mathcal{T}$-blocked.
  3. 3.
    $M_{t},N_{t}$ and $F_{t}$ are $\mu$-stochastic.
Proof.
  1. 1.
    This is clear from the definitions.
  2. 2.
    Since $\tilde{M_{t}}$ is a matching of vertices in $A_{t}^{l}\setminus S_{t}\subseteq A_{t}$ to vertices in $A_{t}^{r}\setminus S_{t}\subseteq A_{t}$, then $\tilde{M_{t}}(i,j)=0$ for all $i\neq j$ such that $(i,j)\notin A_{t}\times A_{t}$. After the addition of $\operatorname{\mathbf{diag}}(\mu-\tilde{M_{t}}\mathbbm{1}_{n})$, we still have $M_{t}(i,j)=0$ for all $i\neq j$ such that $(i,j)\notin A_{t}\times A_{t}$, so $M_{t}$ is $A_{t}$-blocked. Now, for $N_{t}$, note that for all $i\neq j$ such that $(i,j)\notin A_{t}\times A_{t}$, $N_{t}(i,j)=\frac{\delta-1}{\delta}U(i,j)+\frac{1}{\delta}M_{t}(i,j)=\frac{\delta-1}{\delta}\cdot 0+\frac{1}{\delta}\cdot 0=0$.
    Finally, $F_{t}$ is $\mathcal{T}$-blocked because it a product of $\mathcal{T}$-blocked matrices.
  3. 3.
    For $M_{t}$ this is true because we obtain it by explicitly making $\tilde{M_{t}}$ $\mu$-stochastic. For $N_{t}$, note that
    $N_{t}\mathbbm{1}_{n}=\left(\frac{\delta-1}{\delta}U+\frac{1}{\delta}M_{t}\right)\mathbbm{1}_{n}=\frac{\delta-1}{\delta}\mu+\frac{1}{\delta}\mu=\mu$
    and $\mathbbm{1}_{n}^{\prime}N_{t}=\mu^{\prime}$ follows from symmetry. For $F_{t}$, we use induction on $t$. $F_{0}=U=\operatorname{\mathbf{diag}}(\mu)$ is clearly $\mu$-stochastic. After step $t$,
    $F_{{t+1}}\mathbbm{1}_{n}=N_{t}U^{-1}F_{t}U^{-1}N_{t}\cdot\mathbbm{1}_{n}=N_{t}U^{-1}F_{t}U^{-1}\cdot\mu=N_{t}U^{-1}F_{t}\mathbbm{1}_{\mathcal{T}}\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu4\mkern-1.5mu)}}{{=}}N_{t}U^{-1}\cdot\mu=N_{t}\mathbbm{1}_{\mathcal{T}}\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu6\mkern-1.5mu)}}{{=}}\mu,$
    where Equalities $(4)$ and $(6)$ follow from Lemma 4.2 $(2)$. We get $\mathbbm{1}_{n}^{\prime}F_{{t+1}}=\mu^{\prime}$ from symmetry.
The following lemmas show that $F_{t}$ is embeddable in $G_{t}$ and that $G_{t}$ is embeddable in $G$. The proof of Lemma 4.3 is deferred to Appendix A.
Lemma 4.3.
Let $F\in\mathop{\mathbb{R}}^{n\times n}$ be a $\mu$-stochastic matrix, and let $H$ be a graph. Fix any $t$. Assume that $F$ is embeddable in $H$ such that the congestion on each edge $e\in H$ is $c(e)$, then
  1. 1.
    $N_{t}\cdot U^{-1}F$ is embeddable in $H\cup M_{t}$ such that the congestion on edges of $M_{t}$ is $\frac{2}{\delta}$ and the congestion on each edge $e\in H$ is still $c(e)$.
  2. 2.
    $FU^{-1}\cdot N_{t}$ is embeddable in $H\cup M_{t}$ such that the congestion on edges of $M_{t}$ is $\frac{2}{\delta}$ and the congestion on each edge $e\in H$ is still $c(e)$.
  3. 3.
    $N_{t}\cdot U^{-1}F_{t}U^{-1}N_{t}$ is embeddable in $H\cup M_{t}$ such that the congestion on edges of $M_{t}$ is $\frac{4}{\delta}$ and the congestion on each edge $e\in H$ is still $c(e)$.
Lemma 4.4.
For all rounds $t$, $F_{t}$ is embeddable in $G_{t}$ with congestion $\frac{4}{\delta}$.
Proof.
This is a direct consequence of Lemma 4.3. ∎
Lemma 4.5.
For all rounds $t$, $G_{t}$ is embeddable in $G$ with congestion $ct$.
Proof.
For every $t$, by the definition of the flow problem at round $t$, $M_{t}$ is embeddable in $G$ with congestion $c$. Summing these routings gives a routing of $G_{t}=\bigcup_{i=1}^{t}{M_{i}}$ in $G$ with congestion $c\cdot t$. ∎
4.2 $A_{T}$ is a near expander in $F_{T}$
In this section we prove that after $T=\Theta(\log^{2}n)$ rounds, with high probability, $A_{T}$ is a near $(\Omega(1),\mu)$-expander in $F_{T}$, which will imply that it is a near $(\Omega(\phi),\mu)$-expander in $G$.
The section is organized as follows. Lemma 4.6 proves matrix identities and Lemma 4.7 specifies a spectral property that our proof requires. We then define a potential function and lower bound the decrease in potential in Lemmas 4.8-4.11. Finally, in Lemma 4.12 and Corollary 4.13 we use the upper bound on the potential at round $T$, to show that with high probability $A_{T}$ is a near $(\Omega(1),\mu)$-expander in $F_{T}$ and a near $(\Omega(\phi),\mu)$-expander in $G$.
Lemma 4.6.
The following relations hold for all $t$:
  1. 1.
    For any $A_{t}\cap\mathcal{T}$-blocked $\mu_{t}$-stochastic matrix $B\in\mathop{\mathbb{R}}^{n\times n}$ we have $I_{t}U^{-\frac{1}{2}}BU^{-\frac{1}{2}}=U^{-\frac{1}{2}}BU^{-\frac{1}{2}}I_{t}$ and $P_{t}\cdot U^{-\frac{1}{2}}BU^{-\frac{1}{2}}=U^{-\frac{1}{2}}BU^{-\frac{1}{2}}\cdot P_{t}$.
  2. 2.
    $I_{t}P_{t}=P_{t}$, $I_{t}^{2}=I_{t}$ and $P_{t}^{2}=P_{t}$.
  3. 3.
    $P_{t}P_{{t+1}}=P_{{t+1}}P_{t}=P_{{t+1}}$.
  4. 4.
    $P_{t}=U^{-\frac{1}{2}}\mathcal{L}(\frac{1}{\mu(A_{t})}\mu_{t}\mu_{t}^{\prime})U^{-\frac{1}{2}}.$
  5. 5.
    for any $v\in\mathop{\mathbb{R}}^{n}$, it holds that $v^{\prime}\mathcal{L}\left(\frac{1}{\mu(A_{t})}\mu_{t}\mu_{t}^{\prime}\right)v=\left\|U_{t}^{\frac{1}{2}}v\right\|_{2}^{2}-\frac{1}{\mu(A_{t})}\left\langle v,\mu_{t}\right\rangle^{2}$.
  6. 6.
    For any $B\in\mathop{\mathbb{R}}^{n\times n}$, $\operatorname{Tr}(I_{t}BB^{\prime})=\sum_{i\in A_{t}\cap\mathcal{T}}\left\|B(i)\right\|_{2}^{2}$.
Proof.
  1. 1.
    Since $B$ is a $A_{t}\cap\mathcal{T}$-blocked $\mu_{t}$-stochastic matrix, $I_{t}\cdot U^{-\frac{1}{2}}BU^{-\frac{1}{2}}=U^{-\frac{1}{2}}BU^{-\frac{1}{2}}\cdot I_{t}$ is clear. Moreover
    $U^{-\frac{1}{2}}BU^{-\frac{1}{2}}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}=U^{-\frac{1}{2}}B\mathbbm{1}_{t}\sqrt{\mu_{t}^{\prime}}=U^{-\frac{1}{2}}\mu_{t}\sqrt{\mu_{t}^{\prime}}=\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}},$
    $\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}U^{-\frac{1}{2}}BU^{-\frac{1}{2}}=\sqrt{\mu_{t}}\mathbbm{1}_{t}^{\prime}BU^{-\frac{1}{2}}=\sqrt{\mu_{t}}\mu_{t}^{\prime}U^{-\frac{1}{2}}=\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}\ .$
    Thus,
    $P_{t}\cdot U^{-\frac{1}{2}}BU^{-\frac{1}{2}}$ $=(I_{t}-\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}})U^{-\frac{1}{2}}BU^{-\frac{1}{2}}$
    $=U^{-\frac{1}{2}}BU^{-\frac{1}{2}}I_{t}-\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}=U^{-\frac{1}{2}}BU^{-\frac{1}{2}}\cdot P_{t}\ .$
  2. 2.
    The first and second equalities are clear from the definitions. For $P_{t}$,
    $P_{t}^{2}$ $=\left(I_{t}-\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}\right)\cdot\left(I_{t}-\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}\right)$
    $=I_{t}-\frac{2}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}+\frac{1}{\mu(A_{t})^{2}}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}=I_{t}-\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}=P_{t}$
  3. 3.
    We show $P_{t}P_{{t+1}}=P_{{t+1}}$. The other direction follows by symmetry.
    $P_{t}P_{{t+1}}$ $=\left(I_{t}-\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}\right)\cdot\left(I_{{t+1}}-\frac{1}{\mu(A_{t+1})}\sqrt{\mu_{t+1}}\sqrt{\mu_{t+1}^{\prime}}\right)$
    $=I_{{t+1}}-\frac{1}{\mu(A_{t+1})}\sqrt{\mu_{t+1}}\sqrt{\mu_{t+1}^{\prime}}-\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}I_{{t+1}}+\frac{1}{\mu(A_{t})\cdot\mu(A_{t+1})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}}\sqrt{\mu_{t+1}}\sqrt{\mu_{t+1}^{\prime}}$
    $=I_{{t+1}}-\frac{1}{\mu(A_{t+1})}\sqrt{\mu_{t+1}}\sqrt{\mu_{t+1}^{\prime}}-\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t+1}^{\prime}}+\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t+1}^{\prime}}$
    $=I_{{t+1}}-\frac{1}{\mu(A_{t+1})}\sqrt{\mu_{t+1}}\sqrt{\mu_{t+1}^{\prime}}=P_{{t+1}}.$
  4. 4.
    Since $\langle\mu_{t},\mathbbm{1}\rangle=\mu(A_{t})$, we get that the degree matrix of $\frac{1}{\mu(A_{t})}\mu_{t}\mu_{t}^{\prime}$ is $U_{t}$. Hence
    $\mathcal{L}\left(\frac{1}{\mu(A_{t})}\mu_{t}\mu_{t}^{\prime}\right)=U_{t}-\frac{1}{\mu(A_{t})}\mu_{t}\mu_{t}^{\prime}=U^{\frac{1}{2}}(I_{t}-\frac{1}{\mu(A_{t})}\sqrt{\mu_{t}}\sqrt{\mu_{t}^{\prime}})U^{\frac{1}{2}}=U^{\frac{1}{2}}P_{t}U^{\frac{1}{2}}.$
    The result follows since $P_{t}$ is $A_{t}\cap\mathcal{T}$-blocked (and has zeros on the diagonal outside of the block).
  5. 5.
    $v^{\prime}\mathcal{L}\left(\frac{1}{\mu(A_{t})}\mu_{t}\mu_{t}^{\prime}\right)v$ $=v^{\prime}(U_{t}-\frac{1}{\mu(A_{t})}\mu_{t}\mu_{t}^{\prime})v=v^{\prime}U_{t}v-\frac{1}{\mu(A_{t})}v^{\prime}\mu_{t}\mu_{t}^{\prime}v$
    $=(U_{t}^{\frac{1}{2}}v)^{\prime}(U_{t}^{\frac{1}{2}}v)-\frac{1}{\mu(A_{t})}(v^{\prime}\mu_{t})(\mu_{t}^{\prime}v)=\left\|U_{t}^{\frac{1}{2}}v\right\|_{2}^{2}-\frac{1}{\mu(A_{t})}\langle v,\mu_{t}\rangle^{2}.$
  6. 6.
    Observe that $\sum_{i\in A_{t}\cap\mathcal{T}}\left\|B(i)\right\|_{2}^{2}=\left\|I_{t}B\right\|_{F}^{2}$. Indeed, $X=I_{t}B$ satisfies $X(i,j)=B(i,j)$ if $i\in A_{t}\cap\mathcal{T}$ (and otherwise $X(i,j)=0$). Therefore
    $\sum_{i\in A_{t}\cap\mathcal{T}}\left\|B(i)\right\|_{2}^{2}$ $=\left\|I_{t}B\right\|_{F}^{2}=\operatorname{Tr}((I_{t}B)(I_{t}B)^{\prime})=\operatorname{Tr}(I_{t}BB^{\prime}I_{t})$
    $=\operatorname{Tr}(I_{t}^{2}BB^{\prime})=\operatorname{Tr}(I_{t}BB^{\prime}).$
    Where the fourth equality follows from Fact B.1 and the last equality follows from (2).
We define the potential $\psi(t)=\operatorname{Tr}[W_{t}^{2}]=\sum_{i\in A_{t}}{\left\|W_{t}(i)\right\|_{2}^{2}}$, where $W_{t}$ was defined as $W_{t}=(P_{t}U^{-\frac{1}{2}}F_{t}U^{-\frac{1}{2}}P_{t})^{\delta}$. Intuitively, by projecting using $P_{t}$ , the potential only “cares” about the vertices of $A_{t}\cap\mathcal{T}$. We prove in Lemma 4.12 that having small potential will certify that $A_{T}$ is a near expander in $F_{t}$.
Before we bound the decrease in potential, we recall Definition 2.4 of a normalized Laplacian $\mathcal{N}(A)=U^{-\frac{1}{2}}\mathcal{L}(A)U^{-\frac{1}{2}}=I_{\mathcal{T}}-U^{-\frac{1}{2}}AU^{-\frac{1}{2}}$, where $A$ is a symmetric $\mathcal{T}$-blocked $\mu$-stochastic matrix. The proof of the following technical lemma appears in Appendix A.
Lemma 4.7.
For any matrix $A\in\mathop{\mathbb{R}}^{n\times n}$, $\operatorname{Tr}(A^{\prime}(I_{\mathcal{T}}-(U^{-\frac{1}{2}}N_{t}U^{-\frac{1}{2}})^{4\delta})A)\geq\frac{1}{3}\operatorname{Tr}(A^{\prime}\mathcal{N}(M_{t})A)$.
In the following lemma we bound the decrease in potential. The bound consists of the contribution of the matched vertices and it also takes into account the removal of $S_{t}$ from $A_{t}$.
Lemma 4.8.
For each round $t$,
$\psi(t)-\psi(t+1)\geq\frac{1}{3}\sum_{\{i,k\}\in M_{t}}w_{ik}\left\|\left(\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right)\right\|_{2}^{2}+\sum_{j\in S_{t}\cap\mathcal{T}}\mu(j)\left\|\frac{W_{t}(j)}{\sqrt{\mu(j)}}\right\|_{2}^{2}$
Proof.
To simplify the notation, we denote $\bar{N}_{t}\mathrel{\mathop{:}}=U^{-\frac{1}{2}}N_{t}U^{-\frac{1}{2}}$ and $\bar{F}_{t}\mathrel{\mathop{:}}=U^{-\frac{1}{2}}F_{t}U^{-\frac{1}{2}}$. We rewrite the potential in the next iteration as follows:
$\psi(t+1)$ $=\operatorname{Tr}(W_{{t+1}}^{2})=\operatorname{Tr}\left(\left(P_{{t+1}}U^{-\frac{1}{2}}F_{{t+1}}U^{-\frac{1}{2}}P_{{t+1}}\right)^{2\delta}\right)$
$=\operatorname{Tr}\left(\left(P_{{t+1}}U^{-\frac{1}{2}}(N_{{t}}U^{-1}F_{{t}}U^{-1}N_{{t}})U^{-\frac{1}{2}}P_{{t+1}}\right)^{2\delta}\right)$
$=\operatorname{Tr}\left(\left(P_{{t+1}}U^{-\frac{1}{2}}(N_{{t}}U^{-\frac{1}{2}}U^{-\frac{1}{2}}F_{{t}}U^{-\frac{1}{2}}U^{-\frac{1}{2}}N_{{t}})U^{-\frac{1}{2}}P_{{t+1}}\right)^{2\delta}\right)$
$=\operatorname{Tr}\left(\left(P_{{t+1}}\bar{N}_{t}\bar{F}_{t}\bar{N}_{t}P_{{t+1}}\right)^{2\delta}\right)$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu6\mkern-1.5mu)}}{{=}}\operatorname{Tr}\left(\left(\bar{N}_{t}P_{{t+1}}\bar{F}_{t}P_{{t+1}}\bar{N}_{t}\right)^{2\delta}\right)$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu7\mkern-1.5mu)}}{{=}}\operatorname{Tr}\left(\left(\bar{N}_{t}P_{{t+1}}P_{t}\bar{F}_{t}P_{t}P_{{t+1}}\bar{N}_{t}\right)^{2\delta}\right)$
$=\operatorname{Tr}\left(\left(\bar{N}_{t}P_{{t+1}}(P_{t}\bar{F}_{t}P_{t})P_{{t+1}}\bar{N}_{t}\right)^{2\delta}\right)\ ,$
where equality $(6)$ follows from Lemma 4.6 (1) for $N_{{t}}$ (which is $A_{t+1}$-blocked $\mu_{t+1}$-stochastic by Lemma 4.2), and equality $(7)$ follows from Lemma 4.6 (3).
By Properties (1) and (2) of Lemma 4.6 it holds that $\bar{N}_{t+1}P_{{t+1}}=P_{{t+1}}\bar{N}_{t+1}=P_{{t+1}}\bar{N}_{t+1}P_{{t+1}}$. Therefore, the potential can be written in terms of symmetric matrices:
$\psi(t+1)$ $=\operatorname{Tr}\left(\left((P_{{t+1}}\bar{N}_{t}P_{{t+1}})(P_{t}\bar{F}_{t}P_{t})(P_{{t+1}}\bar{N}_{t}P_{{t+1}})\right)^{2\delta}\right)$
$\leq\operatorname{Tr}((P_{{t+1}}\bar{N}_{t}P_{{t+1}})^{2\delta}(P_{t}\bar{F}_{t}P_{t})^{2\delta}(P_{{t+1}}\bar{N}_{t}P_{{t+1}})^{2\delta})$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{=}}\operatorname{Tr}((P_{{t+1}}\bar{N}_{t}P_{{t+1}})^{4\delta}(P_{t}\bar{F}_{t}P_{t})^{2\delta})=\operatorname{Tr}((\bar{N}_{t}P_{{t+1}})^{4\delta}W_{t}^{2})$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu4\mkern-1.5mu)}}{{=}}\operatorname{Tr}(\bar{N}_{t}^{4\delta}P_{{t+1}}W_{t}^{2})\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu5\mkern-1.5mu)}}{{=}}\operatorname{Tr}(\bar{N}_{t}^{2\delta}P_{{t+1}}\bar{N}_{t}^{2\delta}W_{t}^{2})$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu6\mkern-1.5mu)}}{{=}}\operatorname{Tr}(W_{t}\bar{N}_{t}^{2\delta}P_{{t+1}}\bar{N}_{t}^{2\delta}W_{t})$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu7\mkern-1.5mu)}}{{=}}\operatorname{Tr}(W_{t}\bar{N}_{t}^{2\delta}U^{-\frac{1}{2}}\mathcal{L}\left(\frac{1}{\mu(A_{t+1})}\mu_{t+1}\mu_{t+1}^{\prime}\right)U^{-\frac{1}{2}}\bar{N}_{t}^{2\delta}W_{t})$
$=\operatorname{Tr}\left(\left(U^{-\frac{1}{2}}\cdot\bar{N}_{t}^{2\delta}W_{t}\right)^{\prime}\cdot\mathcal{L}\left(\frac{1}{\mu(A_{t+1})}\mu_{t+1}\mu_{t+1}^{\prime}\right)\cdot\left(U^{-\frac{1}{2}}\cdot\bar{N}_{t}^{2\delta}W_{t}\right)\right)\ ,$
where the inequality follows from Theorem B.3, equality $(2)$ follows from Fact B.1, equalities $(4)$ and $(5)$ follow from Properties (1) and (2) of Lemma 4.6 (and from the fact that $N_{{t}}$ is $A_{t+1}$-blocked $\mu_{t+1}$-stochastic, by Lemma 4.2), equality $(6)$ again uses Fact B.1 and equality $(7)$ follows from Lemma 4.6 (4).
Let $Z_{t}=U^{-\frac{1}{2}}\cdot\bar{N}_{t}^{2\delta}W_{t}$. By applying Lemma 4.6 (5) we get
$\psi(t+1)$ $\leq\operatorname{Tr}\left(Z_{t}^{\prime}\mathcal{L}\left(\frac{1}{\mu(A_{t+1})}\mu_{t+1}\mu_{t+1}^{\prime}\right)Z_{t}\right)=\sum_{i=1}^{n}(Z_{t}(,i))^{\prime}\mathcal{L}\left(\frac{1}{\mu(A_{t+1})}\mu_{t+1}\mu_{t+1}^{\prime}\right)Z_{t}(,i)$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{=}}\sum_{i=1}^{n}\left(\left\|U_{{t+1}}^{\frac{1}{2}}Z_{t}(,i)\right\|_{2}^{2}-\frac{1}{\mu(A_{t+1})}\left\langle Z_{t}(,i),\mu_{t+1}\right\rangle^{2}\right)\leq\sum_{i=1}^{n}\left\|U_{{t+1}}^{\frac{1}{2}}Z_{t}(,i)\right\|_{2}^{2}$
$=\sum_{i=1}^{n}\sum_{j\in A_{t+1}}\left(\sqrt{\mu(j)}Z_{t}(j,i)\right)^{2}=\sum_{j\in A_{t+1}}\left\|\left(U_{{t+1}}^{\frac{1}{2}}Z_{t}\right)(j)\right\|_{2}^{2}\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu5\mkern-1.5mu)}}{{=}}\sum_{j\in A_{t+1}}\left\|\left(\bar{N}_{t}^{2\delta}W_{t}\right)(j)\right\|_{2}^{2}$
$=\sum_{j\in A_{t}}\left\|\left(\bar{N}_{t}^{2\delta}W_{t}\right)(j)\right\|_{2}^{2}-\sum_{j\in S_{t}}\left\|\left(\bar{N}_{t}^{2\delta}W_{t}\right)(j)\right\|_{2}^{2},$ (1)
where equality $(2)$ holds by Property (5) of Lemma 4.6 and equality $(5)$ holds since we only sum rows in $A_{t+1}$. Since $\bar{N}_{t}$ is diagonal outside $A_{t+1}$ (By the definition of $M_{t}$), we have that $\left(\bar{N}_{t}^{2\delta}W_{t}\right)(j)=W_{t}(j)$, for every $j\in S_{t}$. Thus,
$\sum_{j\in S_{t}}\left\|\left(\bar{N}_{t}^{2\delta}W_{t}\right)(j)\right\|_{2}^{2}=\sum_{j\in S_{t}}\left\|W_{t}(j)\right\|_{2}^{2}.$ (2)
By Lemma 4.6 (6), we get
$\sum_{j\in A_{t}}\left\|\left(\bar{N}_{t}^{2\delta}W_{t}\right)(j)\right\|_{2}^{2}$ $=\sum_{j\in A_{t}\cap\mathcal{T}}\left\|\left(\bar{N}_{t}^{2\delta}W_{t}\right)(j)\right\|_{2}^{2}$ (3)
$=\operatorname{Tr}(I_{{t}}\cdot\bar{N}_{t}^{2\delta}\cdot W_{t}^{2}\cdot\bar{N}_{t}^{2\delta})$
$=\operatorname{Tr}(\bar{N}_{t}^{2\delta}\cdot I_{{t}}\cdot W_{t}^{2}\cdot\bar{N}_{t}^{2\delta})$
$=\operatorname{Tr}(\bar{N}_{t}^{2\delta}\cdot W_{t}^{2}\cdot\bar{N}_{t}^{2\delta})$
$=\operatorname{Tr}(\bar{N}_{t}^{4\delta}W_{t}^{2})$ (4)
where the first equality holds because $N_{t}(i)=0$ for $i\in T$. The second equality holds by Lemma 4.6 (6). The third equality holds since $N_{{t}}$ is $A_{t+1}\cap\mathcal{T}$-blocked $\mu_{t+1}$-stochastic (by Lemma 4.2), so in particular it is $A_{t}\cap\mathcal{T}$-blocked $\mu_{t}$-stochastic, and we can use Lemma 4.6 (1). The fourth equality holds because $I_{t}W_{t}=I_{t}(P_{t}\bar{F}_{t}P_{t})^{\delta}$ and $I_{t}P_{t}=P_{t}$ (by Lemma 4.6 (2)), and the last equality follows from Fact B.1. Plugging Equations (2) and (4) into (1) we get the following bound on the decrease in potential:
$\psi(t)-\psi(t+1)$ $\geq\operatorname{Tr}((I-\bar{N}_{t}^{4\delta})W_{t}^{2})+\sum_{j\in S_{t}}\left\|W_{t}(j)\right\|_{2}^{2}$
$=\operatorname{Tr}(W_{t}(I_{\mathcal{T}}-\bar{N}_{t}^{4\delta})W_{t})+\sum_{j\in S_{t}}\left\|W_{t}(j)\right\|_{2}^{2}$
$\geq\frac{1}{3}\operatorname{Tr}(W_{t}\mathcal{N}(M_{{t}})W_{t})+\sum_{j\in S_{t}}\left\|W_{t}(j)\right\|_{2}^{2}$
$=\frac{1}{3}\operatorname{Tr}((U^{-\frac{1}{2}}W_{t})^{\prime}\mathcal{L}(M_{{t}})(U^{-\frac{1}{2}}W_{t}))+\sum_{j\in S_{t}\cap\mathcal{T}}\mu(j)\left\|\frac{W_{t}(j)}{\sqrt{\mu(j)}}\right\|_{2}^{2}$
$=\frac{1}{3}\sum_{\{i,k\}\in M_{t}}{w_{ik}\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right\|_{2}^{2}}+\sum_{j\in S_{t}\cap\mathcal{T}}\mu(j)\left\|\frac{W_{t}(j)}{\sqrt{\mu(j)}}\right\|_{2}^{2}$
where the first equality follows from Fact B.1 and since as implied before $I_{\mathcal{T}}W_{t}=W_{t}$, the second inequality follows Lemma 4.7, and the last equality follows from Lemma B.5. ∎
The following shows that the potential is expected to drop by a factor of $1-\Omega(1/\log n)$.
Lemma 4.9.
For each round $t$,
$\mathop{\mathbb{E}}\left[\frac{1}{3}\sum_{\{i,k\}\in M_{t}}w_{ik}\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right\|_{2}^{2}+\sum_{j\in S_{t}\cap\mathcal{T}}\mu(j)\left\|\frac{W_{t}(j)}{\sqrt{\mu(j)}}\right\|_{2}^{2}\right]\geq\frac{1}{3000\alpha\log n}\psi(t)-\frac{3}{n^{\alpha/16}}$
for some constant $\alpha$, where the expectation is taken over the unit vector $r\in\mathop{\mathbb{R}}^{n}$.
Proof.
Recall that $u_{i}=\frac{1}{\sqrt{\mu(i)}}\langle W_{t}(i),r\rangle$ for $i\in A_{t}\cap\mathcal{T}$. Notice that $\left(\frac{W_{t}(i)}{\sqrt{\mu(i)}}\right)(j)=\frac{W_{t}(i,j)}{\sqrt{\mu(i)}}$. Use Lemma C.3 from Appendix C for the set of vectors $\left\{\frac{1}{\sqrt{\mu(i)}}W_{t}(i)\mid i\in A_{t}\cap\mathcal{T}\right\}$. By Lemma C.3(2), we have with high probability:
$\forall i,k\in A_{t}\cap\mathcal{T}$ $:\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right\|_{2}^{2}\geq\frac{n}{\alpha\log n}\cdot(u_{i}-u_{k})^{2}$
$\forall i\in A_{t}\cap\mathcal{T}$ $:\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}\right\|_{2}^{2}\geq\frac{n}{\alpha\log n}\cdot u_{i}^{2}$
for some constant $\alpha>0$. In order to replace the inequality with high probability by an inequality in expected values, we introduce a random variable $z$ that is non-zero only when this inequality fails to hold, such that
$\forall i,k\in A_{t}\cap\mathcal{T}$ $:\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right\|_{2}^{2}\geq\frac{n}{\alpha\log n}\cdot(u_{i}-u_{k})^{2}-z$
$\forall i\in A_{t}\cap\mathcal{T}$ $:\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}\right\|_{2}^{2}\geq\frac{n}{\alpha\log n}\cdot u_{i}^{2}-z$
holds with probability $1$. I.e., we define
$\mathcal{B}=\{0\}$ $\cup\left\{\frac{n}{\alpha\log n}(u_{i}-u_{k})^{2}-\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right\|_{2}^{2}:(i,k)\in(A_{t}\cap\mathcal{T})\times(A_{t}\cap\mathcal{T})\right\}$
$\cup\left\{\frac{n}{\alpha\log n}u_{i}^{2}-\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}\right\|_{2}^{2}:i\in A_{t}\cap\mathcal{T}\right\}$
and $z=\max(\mathcal{B})$. Let $K\coloneqq\langle\mathbbm{1},\mu\rangle=poly(m)$. We get that
$\frac{1}{3}\sum_{\{i,k\}\in M_{t}}w_{ik}\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right\|_{2}^{2}$ $\geq\frac{n}{3\alpha\log{n}}\sum_{\{i,k\}\in M_{t}}w_{ik}(u_{i}-u_{k})^{2}-K\cdot z$
$\sum_{j\in S_{t}}\mu(j)\left\|\frac{W_{t}(j)}{\sqrt{\mu(j)}}\right\|_{2}^{2}$ $\geq\frac{n}{\alpha\log{n}}\sum_{j\in S_{t}\cap\mathcal{T}}\mu(j)u_{j}^{2}-K\cdot z.$
This means that
$\frac{1}{3}\sum_{\{i,k\}\in M_{t}}w_{ik}\left\|\left(\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right)\right\|_{2}^{2}+\sum_{j\in S_{t}\cap\mathcal{T}}\mu(j)\left\|\frac{W_{t}(j)}{\sqrt{\mu(j)}}\right\|_{2}^{2}$
$\geq\frac{n}{3\alpha\log{n}}\sum_{\{i,k\}\in M_{t}}w_{ik}(u_{i}-u_{k})^{2}+\frac{n}{\alpha\log{n}}\sum_{j\in S_{t}\cap\mathcal{T}}\mu(j)u_{j}^{2}-2K\cdot z$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu2\mkern-1.5mu)}}{{\geq}}\frac{n}{3\alpha\log{n}}\sum_{i\in A_{t}^{l}\setminus S_{t}}m_{i}(u_{i}-\eta)^{2}+\frac{n}{\alpha\log{n}}\sum_{j\in S_{t}}\mu(j)u_{j}^{2}-2K\cdot z$
$\geq\frac{n}{3\alpha\log{n}}\sum_{i\in A_{t}^{l}\setminus S_{t}}m_{i}(u_{i}-\eta)^{2}+\frac{n}{\alpha\log{n}}\sum_{j\in A_{t}^{l}\cap S_{t}}\mu(j)u_{j}^{2}-2K\cdot z$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu4\mkern-1.5mu)}}{{\geq}}\frac{n}{27\alpha\log{n}}\sum_{i\in A_{t}^{l}\setminus S_{t}}m_{i}u_{i}^{2}+\frac{n}{\alpha\log{n}}\sum_{j\in A_{t}^{l}\cap S_{t}}\mu(j)u_{j}^{2}-2K\cdot z$
$\stackrel{{\scriptstyle\scriptscriptstyle(\mkern-1.5mu5\mkern-1.5mu)}}{{\geq}}\frac{n}{27\alpha\log{n}}\sum_{i\in A_{t}^{l}}m_{i}u_{i}^{2}-2K\cdot z\geq\frac{n}{3000\alpha\log{n}}\sum_{i\in A_{t}}\mu(i)u_{i}^{2}-2K\cdot z\ .$
Inequality $(2)$ is due to Lemma 3.5(1), the fact that $\mu(v)=0$ for every $v\notin\mathcal{T}$ and because each $i\in A^{l}_{t}\setminus S_{t}$ is matched to $A^{r}_{t}\setminus S_{t}$ with a total weight of $m_{i}$, inequality $(4)$ follows from Lemma 3.5(4), inequality $(5)$ is true because $m_{i}\leq\mu(i)$ for all $i\in A^{l}_{t}$, and the last inequality follows from Lemma 3.5(5). Finally, in expectation,
$\mathop{\mathbb{E}}\left[\frac{1}{3}\sum_{\{i,k\}\in M_{t}}w_{ik}\left\|\left(\frac{W_{t}(i)}{\sqrt{\mu(i)}}-\frac{W_{t}(k)}{\sqrt{\mu(k)}}\right)\right\|_{2}^{2}+\sum_{j\in S_{t}\cap\mathcal{T}}\mu(j)\left\|\frac{W_{t}(j)}{\sqrt{\mu(j)}}\right\|_{2}^{2}\right]$
$\geq\frac{n}{3000\alpha\log{n}}\sum_{i\in A_{t}}\mu(i)\mathop{\mathbb{E}}[u_{i}^{2}]-2K\cdot\mathop{\mathbb{E}}[z]$
$=\frac{1}{3000\alpha\log{n}}\sum_{i\in A_{t}\cap\mathcal{T}}\mu(i)\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}\right\|_{2}^{2}-2K\cdot\mathop{\mathbb{E}}[z]$
$=\frac{1}{3000\alpha\log{n}}\sum_{i\in A_{t}}\left\|W_{t}(i)\right\|_{2}^{2}-2K\cdot\mathop{\mathbb{E}}[z]$
$=\frac{1}{3000\alpha\log{n}}\psi(t)-2K\cdot\mathop{\mathbb{E}}[z]$
where the first equality follows from Lemma C.3 (1).
We note that $z=O(poly(n))$. Indeed, for every $i\in A_{t}\cap\mathcal{T}$
$u_{i}=\left\langle\frac{W_{t}(i)}{\sqrt{\mu(i)}},r\right\rangle\leq\left\|\frac{W_{t}(i)}{\sqrt{\mu(i)}}\right\|_{2}=\frac{\left\|W_{t}(i)\right\|_{2}}{\sqrt{\mu(i)}}\leq\frac{1}{\sqrt{\mu(i)}}=O(poly(n)),$
where the first inequality holds due to Cauchy-Schwartz since $r$ is a unit vector and the last inequality holds since all eigenvalues of $W_{t}$ are in $[0,1]$: By Lemma B.6 we get that $U^{-\frac{1}{2}}F_{t}U^{-\frac{1}{2}}=I_{\mathcal{T}}-\mathcal{N}(F_{t})$ has eigenvalues in [1,1][-1,1]. Since $P_{t}$ is a projection matrix, it has eigenvalues in $[0,1]$. Therefore, $P_{t}U^{-\frac{1}{2}}F_{t}U^{-\frac{1}{2}}P_{t}$ has eigenvalues in [1,1][-1,1]. Finally, $W_{t}=(P_{t}U^{-\frac{1}{2}}F_{t}U^{-\frac{1}{2}}P_{t})^{\delta}$ is PSD since $\delta$ is a power of $2$, so its eigenvalues are in $[0,1]$.
By Lemma C.3, $z$ is non-zero with probability at most $\frac{1}{n^{\alpha/8}}$, so for large enough $\alpha$, we get $2K\mathop{\mathbb{E}}[z]\leq\frac{2Kpoly(n)}{n^{\alpha/8}}\leq\frac{3}{n^{\alpha/16}}$, [^10] The second Inequality explains why Assumption 2.1 requires $\min_{v\in\mathcal{T}}\mu(v)\leq poly(n)$. completing the proof. ∎
The following two corollaries follow by Lemmas 4.8 and 4.9.
Corollary 4.10.
For each round $t$, $\mathop{\mathbb{E}}[\psi(t+1)]\leq\left(1-\frac{1}{3000\alpha\log n}\right)\psi(t)+\frac{3}{n^{\alpha/16}}$, where the expectation is over the unit vector $r\in\mathop{\mathbb{R}}^{n}$.
Corollary 4.11 (Total Potential).
With high probability over the choices of $r$, $\psi(T)\leq\frac{1}{n}$.
Lemma 4.12.
Let $H=(V,\bar{E})$ be a graph on $n$ vertices, such that $F_{T}$ is its weighted adjacency matrix. Assume that $\psi(T)\leq\frac{1}{n}$. Then, $A_{T}$ is a near $(\frac{1}{5},\mu)$-expander with respect to $\mu$ in $H$.
Proof.
Recall that $F_{T}$ is symmetric and $\mu_{T}$-stochastic. Let $k=\mu(A_{T})$. Let $S\subseteq A_{T}$ be a cut, and denote $\mu_{S}\in\mathop{\mathbb{R}}^{n}$ to be the vector where $\mu_{S}(u)=\left\{\begin{array}[]{cl}\mu(u)&\mbox{if $u\in S$,}\\ 0&\mbox{otherwise.}\end{array}\right$. Additionally, denote $\ell=\mu(S)\leq\frac{1}{2}k$. Note that $\left\|\sqrt{\mu_{S}}\right\|_{2}^{2}=\ell$.
Denote by $\bar{\lambda}\geq 0$ the largest singular value of $X_{T}\mathrel{\mathop{:}}=P_{T}U^{-\frac{1}{2}}F_{T}U^{-\frac{1}{2}}P_{T}$ (square root of the largest eigenvalue of $(P_{T}U^{-\frac{1}{2}}F_{T}U^{-\frac{1}{2}}P_{T})^{2}$). Because $\operatorname{Tr}(X_{T}^{2\delta})=\psi(T)\leq\frac{1}{n}$, we have in particular that the largest eigenvalue of $X_{T}^{2\delta}$ is at most $\frac{1}{n}$, so we have $\bar{\lambda}\leq\frac{1}{n^{\frac{1}{\delta}}}$. We choose $\delta=\Theta(\log n)$ [^11] This gives the upper bound on $\delta$. The lower bound comes from Lemma 4.4. such that $\frac{1}{n^{\frac{1}{\delta}}}\leq\frac{1}{20}$, so $\bar{\lambda}\leq\frac{1}{20}$.
In order to prove near-expansion we need to lower bound $|E_{F_{T}}(S,\bar{S})|$. We do so by upper bounding $|E_{F_{T}}(S,S)|=\mathbbm{1}_{S}^{\prime}F_{T}\mathbbm{1}_{S}$. Because $F_{t}$ is $\mathcal{T}$-blocked we get that $\mathbbm{1}_{S}^{\prime}F_{T}\mathbbm{1}_{S}=\mathbbm{1}_{S}^{\prime}(I_{T}F_{T}I_{T})\mathbbm{1}_{S}$. Observe the following relation between $X_{T}$ and $I_{T}F_{T}I_{T}$:
$U^{\frac{1}{2}}X_{T}U^{\frac{1}{2}}$ $=U^{\frac{1}{2}}(P_{T}U^{-\frac{1}{2}}F_{T}U^{-\frac{1}{2}}P_{T})U^{\frac{1}{2}}=U^{\frac{1}{2}}(I_{T}-\frac{1}{k}\sqrt{\mu_{T}}\sqrt{\mu_{T}^{\prime}})U^{-\frac{1}{2}}F_{T}U^{-\frac{1}{2}}(I_{T}-\frac{1}{k}\sqrt{\mu_{T}}\sqrt{\mu_{T}^{\prime}})U^{\frac{1}{2}}$
$=(I_{T}-\frac{1}{k}\mu_{T}\mathbbm{1}_{T}^{\prime})F_{T}(I_{T}-\frac{1}{k}\mathbbm{1}_{T}\mu_{T}^{\prime})=I_{T}F_{T}I_{T}-\frac{1}{k}\mu_{T}\mathbbm{1}_{T}^{\prime}F_{T}I_{T}-\frac{1}{k}I_{T}F_{T}\mathbbm{1}_{T}\mu_{T}^{\prime}+\frac{1}{k^{2}}\mu_{T}\mathbbm{1}_{T}^{\prime}F_{T}\mathbbm{1}_{T}\mu_{T}^{\prime}.$
Rearranging the terms, we get
$I_{T}F_{T}I_{T}=U^{\frac{1}{2}}X_{T}U^{\frac{1}{2}}+\frac{1}{k}\mu_{T}\mathbbm{1}_{T}^{\prime}F_{T}I_{T}+\frac{1}{k}I_{T}F_{T}\mathbbm{1}_{T}\mu_{T}^{\prime}-\frac{1}{k^{2}}\mu_{T}\mathbbm{1}_{T}^{\prime}F_{T}\mathbbm{1}_{T}\mu_{T}^{\prime}\ .$
Therefore
$|E_{F_{T}}(S,S)|=$ $\mathbbm{1}_{S}^{\prime}F_{T}\mathbbm{1}_{S}=\mathbbm{1}_{S}^{\prime}\left(U^{\frac{1}{2}}X_{T}U^{\frac{1}{2}}+\frac{1}{k}\mu_{T}\mathbbm{1}_{T}^{\prime}F_{T}I_{T}+\frac{1}{k}I_{T}F_{T}\mathbbm{1}_{T}\mu_{T}^{\prime}-\frac{1}{k^{2}}\mu_{T}\mathbbm{1}_{T}^{\prime}F_{T}\mathbbm{1}_{T}\mu_{T}^{\prime}\right)\mathbbm{1}_{S}.$
We analyze the summands separately. The first summand can be bounded using $\bar{\lambda}$, the largest singular value of $X_{T}$:
$\mathbbm{1}_{S}^{\prime}U^{\frac{1}{2}}X_{T}U^{\frac{1}{2}}\mathbbm{1}_{S}=\sqrt{\mu_{S}^{\prime}}X\sqrt{\mu_{S}}=\left\langle\sqrt{\mu_{S}},X\sqrt{\mu_{S}}\right\rangle\leq\left\|\sqrt{\mu_{S}}\right\|_{2}\left\|X_{T}\sqrt{\mu_{S}}\right\|_{2}\leq\left\|\sqrt{\mu_{S}}\right\|_{2}^{2}\bar{\lambda}\leq\frac{\ell}{20},$
where the first inequality is the Cauchy-Schwartz inequality. Observe that the second and third summands are equal:
$\frac{1}{k}\mathbbm{1}_{S}^{\prime}\mu_{T}\mathbbm{1}_{T}^{\prime}F_{T}I_{T}\mathbbm{1}_{S}=\frac{\ell}{k}\mathbbm{1}_{T}^{\prime}F_{T}\mathbbm{1}_{S}=\frac{\ell}{k}\mathbbm{1}_{S}^{\prime}F_{T}\mathbbm{1}_{T}=\frac{1}{k}\mathbbm{1}_{S}^{\prime}I_{T}F_{T}\mathbbm{1}_{T}\mu_{T}^{\prime}\mathbbm{1}_{S},$
where the second equality follows by transposing and since $F_{T}$ is symmetric. We now bound the sum of the second, third and fourth summands:
$\mathbbm{1}_{S}^{\prime}\left(\frac{1}{k}\mu_{T}\mathbbm{1}_{T}^{\prime}F_{T}I_{T}+\frac{1}{k}I_{T}F_{T}\mathbbm{1}_{T}\mu_{T}^{\prime}-\frac{1}{k^{2}}\mu_{T}\mathbbm{1}_{T}^{\prime}F_{T}\mathbbm{1}_{T}\mu_{T}^{\prime}\right)\mathbbm{1}_{S}=\frac{2\ell}{k}\mathbbm{1}_{T}^{\prime}F_{T}\mathbbm{1}_{S}-\frac{\ell^{2}}{k^{2}}\mathbbm{1}_{T}^{\prime}F_{T}\mathbbm{1}_{T}$
$\leq\left(\frac{2\ell}{k}-\frac{\ell^{2}}{k^{2}}\right)\mathbbm{1}_{T}^{\prime}F_{T}\mathbbm{1}_{S}\leq\left(\frac{2\ell}{k}-\frac{\ell^{2}}{k^{2}}\right)\mathbbm{1}^{\prime}F_{T}\mathbbm{1}_{S}=\left(\frac{2\ell}{k}-\frac{\ell^{2}}{k^{2}}\right)\mu^{\prime}\mathbbm{1}_{S}=\frac{\ell}{k}\left(2-\frac{\ell}{k}\right)\ell,$
where the first inequality follows since $S\subseteq A_{t}$. Note that $\frac{\ell}{k}\in[0,\frac{1}{2}]$. The last inequality is true because for $\frac{\ell}{k}$ in this range, $\left(\frac{2\ell}{k}-\frac{\ell^{2}}{k^{2}}\right)\geq 0$. Moreover, because $\frac{\ell}{k}\in\left[0,\frac{1}{2}\right]$, we have $\frac{\ell}{k}\left(2-\frac{\ell}{k}\right)\leq\frac{3}{4}$. Therefore, $|E_{F_{T}}(S,S)|\leq\frac{1}{20}\ell+\frac{3}{4}\ell=\frac{4}{5}\ell$, and
$|E(S,\bar{S})|$ $=\sum_{u\in S}{\sum_{v\in\bar{S}}{F_{T}(u,v)}}=\sum_{u\in S}{\sum_{v\in V}{F_{T}(u,v)}}-\sum_{u\in S}{\sum_{v\in S}{F_{T}(u,v)}}$
$=\sum_{u\in S}{\mu(u)}-\sum_{u\in S}{\sum_{v\in S}{F_{T}(u,v)}}\geq\ell-\frac{4}{5}\ell=\frac{\ell}{5}\ .$
So $\Phi_{G}(S,\bar{S})=\frac{|E(S,\bar{S})|}{\mu(S)}\geq\frac{1}{5}$, and this is true for all cuts $S\subseteq A$ with $\frac{\mu(S)}{\mu(A_{t})}\leq\frac{1}{2}$.
Corollary 4.13.
If we reach round $T$, then with high probability, $A_{T}$ is a near $(\Omega(\phi),\mu)$-expander in $G$.
Proof.
Assume we reach round $T$. By Corollary 4.11 and Lemma 4.12, with high probability, $A_{T}$ is a near $(\Omega(1),\mu)$-expander in $F_{T}$. By Lemma 4.4, $F_{T}$ is embeddable in $G_{T}$ with congestion $O(\frac{1}{\delta})$. Therefore, by Lemma 2.9, $A_{T}$ is a near $(\Omega(\delta),\mu)$-expander in $G_{T}$.
Furthermore, by Lemma 4.5, $G_{T}$ is embeddable in $G$ with congestion $cT$. So by Lemma 2.9 again, it follows that $A$ is a near $(\Omega(\frac{\delta}{cT}),\mu)$-expander in $G$. Recall that $c=O\left(\frac{1}{\phi\log n}\right)$, $\delta=\Theta(\log n)$, and $T=O(\log^{2}n)$. Therefore, AA is an near $(\Omega(\phi),\mu)$-expander in $G$. ∎
4.3 Proof of Theorem 3.1
We are now ready to prove Theorem 3.1.
Proof of Theorem 3.1.
Recall that $S_{t}$ denotes the cut returned by Lemma 3.6 (applied on $G[A_{t}]$) at iteration $t$, so that $A_{t+1}=A_{t}\setminus S_{t}$.
Observe first that in any round $t$, we have $\Phi^{\mu}_{G}(A_{t},R_{t})\leq\frac{7}{c}=O(\phi\log n)$. This is because $R_{t}=\bigcup_{0\leq t^{\prime}<t}{S_{t^{\prime}}}$ and by Lemma 3.6, for each $t^{\prime}$, $\Phi^{\mu}_{G[A_{t^{\prime}}]}(S_{t^{\prime}},\bar{S_{t^{\prime}}})\leq\frac{7}{c}=O(\phi\log n)$.
Assume Algorithm 1 terminates because $\mu(R_{t})>\frac{\mu(V)\cdot c\cdot\phi}{70}=\Omega(\frac{\mu(V)}{\log n})$. We also have, by Lemma 3.6, that $\mu(A_{t})=\Omega(\mu(V))=\Omega(\frac{\mu(V)}{\log n})$. Then $(A_{t},R_{t})$ is a balanced cut where $\Phi_{G}(A_{t},R_{t})=O(\phi\log n)$. We end in Case (2) of Theorem 3.1.
Otherwise, Algorithm 1 terminates at round $T$ and we apply Corollary 4.13. If $R=\emptyset$, then we obtain the first case of Theorem 3.1 because the whole vertex set $V$ is, with high probability, a near $(\Omega(\phi),\mu)$-expander, which means that $G$ is an $(\Omega(\phi),\mu)$-expander. Otherwise, we write $c=\frac{c_{1}}{\phi\log n}$ for some constant $c_{1}$, and $c_{0}\mathrel{\mathop{:}}=\frac{7}{c_{1}}$. We have $\Phi_{G}(A_{T},R_{T})\leq\frac{7}{c}=\frac{7}{c_{1}}\phi\log n=c_{0}\phi\log n$. Additionally, $\mu(R_{t})\leq\frac{\mu(V)\cdot c\cdot\phi}{70}=\frac{\mu(V)\cdot c_{1}}{70\log n}=\frac{\mu(V)}{10c_{0}\log n}$, and, with high probability, $A_{T}$ is a near $(\Omega(\phi),\mu)$-expander in $G$, which means we obtain the third case of Theorem 3.1.
Finally, the running time of Algorithm  1 is $\tilde{O}(m)$: The time of iteration $t$ is the sum of the running times of the following steps:
  1. 1.
    Sample a random unit vector $r\in\mathop{\mathbb{R}}^{n}$.
  2. 2.
    Compute the projections vector $u=U^{-\frac{1}{2}}W_{t}\cdot r$. This takes $O(t\cdot\delta\cdot m)=O(m\cdot t\cdot\log n)$ time since $W_{t}$ is a multiplication of $O(t\cdot\delta)$ matrices, where each matrix either has $O(m)$ non-zero entries or is a projection matrix $P_{t}$.
  3. 3.
    Computing $A^{l}_{t}$ and $A^{r}_{t}$ in time $O(n\log n)$ (Lemma 3.5).
  4. 4.
    Computing the cut $S_{t}$ and the flow $f$ on $G-S_{t}$ in time $\tilde{O}(m)$ (Lemma 3.6).
  5. 5.
    Moving $S_{t}$ from $A_{t}$ to $R_{t+1}$ in time $O(m)$.
  6. 6.
    Constructing $M_{t}$ in time $O(m\log n)$ (using dynamic trees [11]).
This gives a total running time of $\tilde{O}(m)$ for iteration $t$. As $t$ ranges from $1$ to $T=\Theta(\log^{2}n)$, this completes the proof of Theorem 3.1. ∎
5 Expander Decomposition
In this section, we present the standard derivation of an expander decomposition (Theorem 3.2) from Theorem 3.1.
The following lemma, called the trimming step, is the last key component for the expander decomposition algorithm. This is a slight generalization of Theorem 8.2 of [7]. For completeness, we include its proof in Appendix A for completeness.
Theorem 5.1.
Given a graph $G=(V,E)$, $A\subseteq V$, a parameter $\phi>0$, and a weight function $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$ over the vertices, such that $A$ is a near $(\phi,\mu)$-expander in $G$ and $|E(A,\bar{A})|\leq\frac{\phi\cdot\mu(A)}{9}$ The trimming step finds with high probability $A^{\prime}\subseteq A$ in time $\tilde{O}(m)$ such that $\Phi^{\mu}(G[A^{\prime}])\geq\phi/6$. Moreover, $\mu(A^{\prime})\geq\mu(A)-\frac{4|E(A,\bar{A})|}{\phi}$, and $|E(A^{\prime},\bar{A^{\prime}})|\leq 2|E(A,\bar{A})|$.
Using Theorem 5.1 and Theorem 3.1, we prove the following improvement on Theorem 3.1.
Theorem 5.2.
Given a graph $G=(V,E)$ of $m$ edges, a parameter $\phi>0$, and a weight function $\mu:V\to\mathop{\mathbb{R}}_{\geq 0}$ over the vertices, there exists a randomized algorithm which takes $\tilde{O}(m)$ time and must end in one of the following three cases:
  1. 1.
    We certify that $G$ has $\mu$-expansion $\Phi^{\mu}(G)=\Omega(\phi)$, with high probability.
  2. 2.
    We find a cut $(A,\bar{A})$ in $G$ of $\mu$-expansion $\Phi^{\mu}_{G}(A,\bar{A})=O(\phi\log n)$, and $\mu(A),\mu(\bar{A})$ are both $\Omega\left(\frac{\mu(V)}{\log n}\right)$, i.e, we find a relatively balanced low $\mu$-expansion cut.
  3. 3.
    We find a cut $(A,\bar{A})$ with $0<\mu(\bar{A})\leq\frac{\mu(V)}{\log n}$, $\Phi^{\mu}_{G}(A,\bar{A})=O(\phi\log n)$, and with high probability, $G[A]$ is $\Omega(\phi)$-expander with respect to $\mu$.
Proof of Theorem 5.2.
We apply Theorem 3.1. Cases (1) and (2) translate directly to Cases (1) and (2) of Theorem 5.2. If Theorem 3.1 terminates with Case (3), we use Theorem 5.1. We have $\mu(\bar{A})\leq\frac{\mu(V)}{10c_{0}\log n}$, and $\Phi_{G}^{\mu}(A,\bar{A})\leq c_{0}\phi\log n$. This means that $|E(A,\bar{A})|\leq\frac{\mu(V)\cdot\phi}{10}\leq\frac{\mu(A)\cdot\phi}{9}$. The trimming step takes $\tilde{O}\left(m\right)$ time, and we return the cut $(A^{\prime},\bar{A^{\prime}})$. It satisfies that $G[A^{\prime}]$ is a $\Omega(\phi)$ expander with respect to $\mu$. Additionally, $\mu(\bar{A^{\prime}})\leq\mu(\bar{A})+\mu(A)-\mu(A^{\prime})\leq\mu(\bar{A})+\frac{4|E(A,\bar{A})|}{\phi}\leq\frac{\mu(V)}{10c_{0}\log n}+\frac{2\mu(V)}{5}\leq\frac{\mu(V)}{2}$. This means that $\bar{A^{\prime}}$ is still the smaller side of the cut, with respect to $\mu$. By Theorem 5.1, we have that $|E(A^{\prime},\bar{A^{\prime}})|\leq 2|E(A,\bar{A})|$, so $\Phi_{G}^{\mu}(A^{\prime},\bar{A^{\prime}})=\frac{|E(A^{\prime},\bar{A^{\prime}})|}{\mu(\bar{A^{\prime}})}\leq 2\frac{|E(A,\bar{A})|}{\mu(\bar{A})}=2\Phi_{G}^{\mu}(A,\bar{A})=O(\phi\log n)$. If $\mu(\bar{A^{\prime}})\leq\frac{\mu(V)}{\log n}$, we can terminate with Case (3) of the theorem ($\mu(\bar{A^{\prime}})\geq\mu(\bar{A})>0$). Otherwise, $\mu(A^{\prime})$, $\mu(\bar{A^{\prime}})$ are both $\Omega\left(\frac{\mu(V)}{\log n}\right)$, so we can terminate with case (2) of the theorem. ∎
We now show how to get an expander decomposition from Theorem 5.2. The following is the expander decomposition procedure of [12].
Algorithm 2 Expander Decomposition [12]
1:function Decomp($G$, $\phi,\mu$)
2:  Call Cut-Matching($G$, $\phi$) $\triangleright$ See Theorem 5.2
3:  if we certify that $\Phi^{\mu}_{G}\geq\phi$ then return $G$
4:  else if we find a relatively balanced cut $(A,R)$ then
5:   return Decomp($G[A]$, $\phi$, $\mu$) $\cup$ Decomp($G[R]$, $\phi$, $\mu$)
6:  else we find a very unbalanced cut $(A,R)$ $\triangleright$ This cut was trimmed - AA is an expander
7:   return Decomp($G[R]$, $\phi$, $\mu$) $\cup A$
8:  end if
9:end function
Proof of Theorem 3.2.
First, note that by definition, the leaves of the recursion tree give an expander decomposition with $\mu$-expansion of $\Omega(\phi)$.
To bound the running time, note that if we get Case (2) of Theorem 5.2 then both sides of the cut have total $\mu(\cdot)$ measure of at most $\left(1-\Omega\left(\frac{1}{\log n}\right)\right)\cdot\mu(V)$. If we get Case (3), then $\mu(A)=\Omega(\mu(V))$, so the total $\mu(\cdot)$ is reduced by a constant factor (i.e., $\mu(\bar{A})\leq\left(1-\Omega(1)\right)\cdot\mu(V)$). In any case, the total $\mu(\cdot)$ measure of the largest component drops by a factor of at least $1-\Omega\left(\frac{1}{\log n}\right)$ across each level of the recursion, so (as $\mu(V)=O(poly(n))$, by Assumption 2.1) the recursion depth is $O(\log^{2}n)$. [^12] To be precise, the recursion depth is $O(\log n\log\mu^{\star})$, where $\mu^{\star}=\frac{\max(\mu(\cdot))}{\min(\mu(\cdot))}$.
Since the components on one level of the recursion are all disjoint, the total running time on all components of one level of the recursion is $\tilde{O}(m)$. Since the depth of the recursion is $O(\log^{2}n)$ we get that the running time is $\tilde{O}(m)$.
To bound the number of edges between expander clusters, observe that in both Case (2) and Case (3), we always cut a component along a cut of $\mu$-expansion $O(\phi\log n)$. Thus, we can charge the edges on the cut to the $\mu(\cdot)$ of the vertices on the smaller side of the cut (with respect to $\mu$), so each vertex $v$ is charged $O(\phi\mu(v)\log n)$. A vertex can be on the smaller side of a cut at most $O(\log n)$ times, so we can charge each vertex $v$ at most [^13] To be precise, we can charge each vertex at most $O(\phi\mu(v)\log n\log\mu^{\star})$. $O(\phi\mu(v)\log^{2}n)$ to pay for all the edges between the final clusters. This bounds the total number of edges between the expanders to be at most $O(\phi\mu(V)\log^{2}n)$. ∎
Appendix A Omitted proofs
This appendix contains the proofs that were deferred from the main text. Section A.1 gives algebraic and embedding lemmas that support the analysis of the cut-matching game. In Section A.2 we introduce the fair-cut framework of [7] and use it to analyze flow-based routines: first the flow subroutine in the cut-matching game (Lemma 3.6), and then the trimming step (Theorem 5.1).
A.1 Algebraic and Embedding Proofs
Proof of Lemma 3.5.
Let $L=\{i\in A_{t}\mid u_{i}<0\}$ and $R=\{i\in A_{t}\mid u_{i}\geq 0\}$ and assume w.l.o.g. that $\mu(L)\leq\mu(R)$. In particular, $\mu(L)\leq\mu(A_{t})/2\leq\mu(R)$. For a set $B\subseteq A_{t}$, denote $P_{B}=\sum_{i\in B}\mu(i)u_{i}^{2}$. We divide into cases.
Case $P_{L}\geq\frac{1}{20}P_{A_{t}}$: We set $\eta=0$, $A_{t}^{r}=R$, and $\bar{m}_{i}=\mu(i)$ for $i\in A_{t}^{r}$. The construction of $A_{t}^{l}$ is as follows: If $\mu(L)\leq\mu(A_{t})/8$, then we set $A_{t}^{l}=L$, and $m_{i}=\mu(i)$ for $i\in A_{t}^{l}$.
Otherwise, we sort $L$ according to $u_{i}$ and scan the vertices (from smallest to largest) until the accumulative $\mu(\cdot)$ weight of the scanned vertices is at least $\mu(A_{t})/8$. Let $i_{1},\ldots,i_{k}$ be the scanned vertices. We set $A_{t}^{l}=\{i_{1},\ldots,i_{k}\}$, $m_{i_{j}}=\mu(i_{j})$ for $1\leq i_{j}<k$, and $m_{i_{k}}=\mu(A_{t})/8-\mu(A_{t}^{l}\setminus\{i_{k}\})$. It is easy to verify that properties (1)(1)-$(3)$ of the lemma hold. Property $(4)$ holds because of the choice $\eta=0$. Finally, property $(5)$ holds since $P_{L}\geq\frac{1}{20}P_{A_{t}}$ and since $A_{t}^{l}$ contains the vertices with the smallest $u_{i}$ values (of total $\mu(\cdot)$ weight equaling $\mu(A_{t})/8$, while $\mu(L)\leq\mu(A_{t})/2$) in $L$.
Case $P_{L}<\frac{1}{20}P_{A_{t}}$: Therefore $P_{R}>\frac{19}{20}P_{A_{t}}$. By the statement of the lemma, we have that $\sum_{i\in A_{t}}{\mu(i)u_{i}}=0$. Denote by $\Delta$ the total distance from zero, that is $\Delta=\sum_{i\in A_{t}}\mu(i)|u_{i}|$. Observe that $\sum_{i\in L}\mu(i)|u_{i}|=\sum_{i\in R}\mu(i)|u_{i}|=\Delta/2$. Let $M=\mu(A_{t})$. We set:
  • $\eta=4\Delta/M$.
  • $A_{t}^{r}=\{i\in A_{t}\mid u_{i}\leq\eta\}$, and $\bar{m}_{i}=\mu(i)$ for $i\in A_{t}^{r}$.
  • $R^{\prime}=\{i\in A_{t}\mid u_{i}\geq 6\Delta/M\}$
  • $A_{t}^{\ell}$ is the subset of $R^{\prime}$ consisting of the vertices with the largest $u_{i}$ values, chosen so that their total $\mu(\cdot)$ weight satisfies $\mu(A_{t}^{l})=\mu(A_{t})/8$. [^14] This is done similarly to the previous case. Again, $m_{i}=\mu(i)$ for every $i\in A_{t}^{l}$, except for possibly one vertex.
It is easy to see that properties $(1),(2)$ of the lemma hold. Note that $\mu(R\setminus A_{t}^{r})\leq\mu(R)/2$, as otherwise
$\Delta=\sum_{i\in A_{t}}\mu(i)|u_{i}|\geq\sum_{i\in R\setminus A_{t}^{r}}\mu(i)|u_{i}|\geq\sum_{i\in R\setminus A_{t}^{r}}\mu(i)\frac{4\Delta}{M}=\mu(R\setminus A_{t}^{r})\frac{4\Delta}{M}>\frac{2\mu(R)\Delta}{M}\geq\Delta,$
a contradiction. Therefore, $\mu(A_{t}^{r})=\mu(L)+\mu(R\cap A_{t}^{r})\geq(\mu(L)+\mu(R))/2=\mu(A_{t})/2$, proving property $(3)$ of the lemma. To see property $(4)$, let $i\in A_{t}^{l}$. Observe that $u_{i}\geq 6\Delta/M$ and $\eta=4\Delta/M$, so $u_{i}-\eta\geq\frac{1}{3}u_{i}$ and therefore $(u_{i}-\eta)^{2}\geq\frac{1}{9}u_{i}^{2}$. We now prove property $(5)$. We show that $P_{R^{\prime}}\geq\frac{1}{20}P_{A_{t}}$ and then the proof is identical to the previous case. By applying the Cauchy–Schwarz inequality on $a_{i}=\sqrt{\mu(i)}u_{i}$ and $b_{i}=\sqrt{\mu(i)}$ for $i\in L$, we get
$P_{L}=\sum_{i\in L}\mu(i)u_{i}^{2}=\langle\vec{a},\vec{a}\rangle\geq\frac{\langle\vec{a},\vec{b}\rangle^{2}}{\langle\vec{b},\vec{b}\rangle}=\frac{\Delta^{2}}{4\mu(L)}\geq\frac{\Delta^{2}}{4M},$
and therefore
$P_{R\setminus R^{\prime}}=\sum_{i\in R\setminus R^{\prime}}\mu(i)u_{i}^{2}<\frac{6\Delta}{M}\sum_{i\in R\setminus R^{\prime}}\mu(i)u_{i}\leq\frac{6\Delta}{M}\sum_{i\in R}\mu(i)u_{i}=\frac{3\Delta^{2}}{M}\leq 12P_{L}.$
Thus, $P_{R^{\prime}}=P_{R}-P_{R\setminus R^{\prime}}>\frac{19}{20}P_{A_{t}}-12P_{L}\geq\frac{7}{20}P_{A_{t}}\geq\frac{1}{20}P_{A_{t}}$. ∎
Proof of Lemma 4.3.
  1. 1.
    Recall that $N_{t}=\frac{\delta-1}{\delta}U+\frac{1}{\delta}M_{t}$. Let $P:E\times V\to\mathop{\mathbb{R}}_{\geq 0}$ be a routing of $F$ in $H$ (where $P((u,v),w)$ indicates how much of $w$’s commodity goes through the edge $(u,v)\in E$ from $u$ to $v$). For brevity, we denote $w_{uv}\coloneqq M_{t}(u,v)$, for every $u,v\in V$. Note that $N_{t}U^{-1}F$ is the flow matrix obtained by performing a weighted average on the rows of $F$ described by $M_{t}$. Explicitly, for every $v\in V,a\in V$, we have
    $(N_{t}U^{-1}F)(v,a)=\left(\frac{\delta-1}{\delta}F+\frac{1}{\delta}M_{t}U^{-1}F\right)(v,a)=\frac{\delta-1}{\delta}F(v,a)+\frac{1}{\delta}\sum_{u:\{u,v\}\in M_{t}}{\frac{w_{vu}}{\mu(u)}F(u,a)}.$
    The precise construction of the new embedding of $N_{t}\cdot U^{-1}F=\frac{\delta-1}{\delta}F+\frac{1}{\delta}M_{t}U^{-1}F$ is as follows.
    $P^{\prime}\leftarrow\frac{\delta-1}{\delta}\cdot P$.
    for $\{a,b\}\in M_{t}$ do
      $P^{\prime}((a,b),a)\leftarrow P^{\prime}((a,b),a)+\frac{w_{ab}}{\delta}$.
      $P^{\prime}((b,a),b)\leftarrow P^{\prime}((b,a),b)+\frac{w_{ab}}{\delta}$.
    end for
    for $\{a,b\}\in M_{t}$ do
      for $(u,v)\in H$ do
       $P^{\prime}((u,v),a)\leftarrow P^{\prime}((u,v),a)+\frac{w_{ab}}{\delta\mu(b)}P((u,v),b)$.
       $P^{\prime}((u,v),b)\leftarrow P^{\prime}((u,v),b)+\frac{w_{ab}}{\delta\mu(a)}P((u,v),a)$.
      end for
    end for
    The argument that we indeed obtain an embedding of $N_{t}U^{-1}F=\frac{\delta-1}{\delta}F+\frac{1}{\delta}M_{t}U^{-1}F$ is as follows. We think of $P^{\prime}$ in stages. In the first stage, we scale $P$ by $\frac{\delta-1}{\delta}$. This routes $\frac{\delta-1}{\delta}F(v,a)$ units of flow (of $v$’s commodity) from $v$ to $a$ for every $v$ and $a$. After this stage, each vertex $v\in V$ currently sends $\frac{(\delta-1)}{\delta}\mu(v)$ units of its commodity. In the next stage we wish to route an additional $\frac{1}{\delta}\sum_{u:\{u,v\}\in M_{t}}{\frac{w_{vu}}{\mu(u)}F(u,a)}$ units from $v$ to $a$. To this end, we first move $\frac{w_{vu}}{\delta}$ units from $v$’s commodity to each $u$ with $\{u,v\}\in M_{t}$ (this flow is sent through the edge $\{v,u\}\in H\cup M_{t}$ from $v$ to $u$). Now, each vertex $u\in V$ “mixes” the commodities it got from its neighbors and routes the $\frac{\mu(u)}{\delta}$ “new” units according to $P$, as if they were of its own commodity. Thus, $\frac{F(u,a)}{\delta}$ is the total flow sent from $u$ to $a$ when it routes the $\frac{\mu(u)}{\delta}$ “new” units and the “share” of $v$’s commodity from this flow is $\frac{w_{vu}}{\mu(u)}$. It follows that out of the $\frac{w_{vu}}{\delta}$ units $v$ sent to $u$, exactly $\frac{w_{vu}}{\mu(u)}\frac{F(u,a)}{\delta}$ units go to $a$.
    As for the congestion, note that on each edge of the matching $M_{t}$, we route $\frac{w_{ab}}{\delta}$ units of flow in each direction, so the congestion on each such edge is $\frac{2}{\delta}$. For each (directed) arc $(u,v)\in H$ and $a\in V$, we have $P^{\prime}((u,v),a)=\frac{\delta-1}{\delta}\cdot P((u,v),a)+\sum_{b:\{a,b\}\in M_{t}}{\frac{w_{ab}}{\delta\mu(b)}P((u,v),b)}$. Therefore, on each arc $(u,v)\in H$, the congestion is
    $\sum_{a\in V}{P^{\prime}((u,v),a)}$ $=\frac{\delta-1}{\delta}\cdot\sum_{a\in V}{P((u,v),a)}+\sum_{a\in V}{\sum_{b:\{a,b\}\in M_{t}}{\frac{w_{ab}}{\delta\mu(b)}P((u,v),b)}}$
    $=\frac{\delta-1}{\delta}\cdot\sum_{a\in V}{P((u,v),a)}+\frac{1}{\delta}\sum_{b\in V}{P((u,v),b)}$
    $=\frac{\delta-1}{\delta}\cdot c(u,v)+\frac{1}{\delta}c(u,v)=c(u,v)\ .$
  2. 2.
    Let $P$ be the routing of $F$ in $H$. Note that $FU^{-1}N_{t}=\frac{\delta-1}{\delta}F+\frac{1}{\delta}FU^{-1}M_{t}$ is the flow matrix obtained by performing a weighted average on the columns of $F$ matched by $U^{-1}M_{t}$, i.e. we average the flow received by matched vertices.
    We define a routing $P^{\prime}$ of $FU^{-1}N_{t}$ in $H\cup M_{t}$ as follows:
    $P^{\prime}\leftarrow P$.
    for $\{a,b\}\in M_{t}$ do
      for $u\in V$ do
       $P^{\prime}((a,b),u)\leftarrow\frac{w_{ab}}{\delta\mu(a)}\cdot F(u,a)$.
       $P^{\prime}((b,a),u)\leftarrow\frac{w_{ba}}{\delta\mu(b)}\cdot F(u,b)$.
      end for
    end for
    That is, $P^{\prime}$ routes the same as $P$ over edges in $H$. For every $\{a,b\}\in M_{t},u\in V$, we set $P^{\prime}((a,b),u)=\frac{w_{ab}}{\delta\mu(a)}\cdot F(u,a)$, and symmetrically, $P^{\prime}((b,a),u)=\frac{w_{ba}}{\delta\mu(b)}\cdot F(u,b)$. That is, for every $\{a,b\}\in M_{t}$, $a$ (respectively, $b$) sends $\frac{w_{ab}}{\delta}$ units of its received flow, which is a mix of commodities (commodity $u$ has a $\frac{F(u,a)}{\mu(a)}$ share in this mix), to $b$ (respectively, $a$). Note that $\left(FU^{-1}N_{t}\right)(u,a)=\frac{\delta-1}{\delta}F(u,a)+\frac{1}{\delta}\sum_{b:\{a,b\}\in M_{t}}{\frac{w_{ab}}{\mu(b)}F(u,b)}$. Thus, $P^{\prime}$ routes $FU^{-1}N_{t}$ in $H$.
    Note that the congestion of $P^{\prime}$ on each edge $e\in H$ is still $c(e)$, and on each arc $(a,b)\in M_{t}$, we have $\sum_{u\in V}{P^{\prime}((a,b),u)}=\sum_{u\in V}{\frac{w_{ab}}{\delta\mu(u)}\cdot F(u,a)}=\frac{w_{ab}}{\delta}$ (the last equality follows as $F$ is $\mu$-stochastic). Therefore at most $\frac{w_{ab}}{\delta}$ flow was routed in each direction, so the congestion on the edge $\{u,v\}\in M_{t}$ is at most $\frac{2}{\delta}$.
  3. 3.
    Note that if $F$ and $N_{t}$ are $\mu$-stochastic (see Lemma 4.2 (2)) then $N_{t}\cdot U^{-1}F$ is also $\mu$-stochastic, since:
    $N_{t}U^{-1}F\cdot\mathbbm{1}_{n}$ $=N_{t}U^{-1}\mu=\mu$
    $\mathbbm{1}_{n}^{\prime}\cdot N_{t}U^{-1}F$ $=\mu^{\prime}\cdot U^{-1}F=\mu^{\prime}\ .$
    Therefore we can apply Parts (1) and (2) to get the result.
Proof of Lemma 4.7.
Let $X_{t}=U^{-\frac{1}{2}}N_{t}U^{-\frac{1}{2}}$. Therefore
$X_{t}=U^{-\frac{1}{2}}\left(\frac{\delta-1}{\delta}U+\frac{1}{\delta}M_{t}\right)U^{-\frac{1}{2}}=\frac{\delta-1}{\delta}I_{\mathcal{T}}+\frac{1}{\delta}U^{-\frac{1}{2}}M_{t}U^{-\frac{1}{2}}=I_{\mathcal{T}}-\frac{1}{\delta}(I_{\mathcal{T}}-U^{-\frac{1}{2}}M_{t}U^{-\frac{1}{2}})=I_{\mathcal{T}}-\frac{1}{\delta}\mathcal{N}(M_{t}).$
Observe that $\mathcal{N}(M_{t})$ and $I_{\mathcal{T}}-X_{t}^{4\delta}$ have the same eigenvectors. Indeed, since both matrices are $\mathcal{T}$-blocked, their null space contains all vectors whose non-zero coordinates are outside $\mathcal{T}$. Consider an eigenvector $v\in\mathop{\mathbb{R}}^{n}$, whose non-zero coordinates are in $\mathcal{T}$, of $\mathcal{N}(M_{t})$, with an eigenvalue of $\lambda$. Then,
$X_{t}v=\left(I_{\mathcal{T}}-\frac{1}{\delta}\mathcal{N}(M_{t})\right)v=v-\frac{\lambda}{\delta}v=\left(1-\frac{\lambda}{\delta}\right)v\ .$
Therefore, $(I_{\mathcal{T}}-X_{t}^{4\delta})v=\left(1-\left(1-\frac{\lambda}{\delta}\right)^{4\delta}\right)v$. We can see that an eigenvector $v\in\mathop{\mathbb{R}}^{n}$ of $\mathcal{N}(M_{t})$, with eigenvalue $\lambda$, is an eigenvector of $I_{\mathcal{T}}-X_{t}^{4\delta}$ with eigenvalue $1-(1-\frac{\lambda}{\delta})^{4\delta}$. A known property of normalized Laplacians (see Lemma B.6) is that all of their eigenvalues are in the interval $[0,2]$, so $\lambda\in[0,2]$. For any such $\lambda$ it holds that $1-(1-\frac{\lambda}{\delta})^{4\delta}\geq 1-\frac{1}{e^{4\lambda}}$. Indeed, for $\lambda=0$ an equality is achieved, and for $\lambda\in(0,2]$, note that $\delta=\Theta(\log n)$, so $\frac{\delta}{\lambda}\geq 1$ and we get
$1-\left(1-\frac{\lambda}{\delta}\right)^{4\delta}=1-\left(\left(1-\frac{\lambda}{\delta}\right)^{\frac{\delta}{\lambda}}\right)^{4\lambda}\geq 1-\frac{1}{e^{4\lambda}}$
Simple calculus shows that $1-\frac{1}{e^{4x}}\geq\frac{1}{3}x$ for any $x\in[0,2]$. Therefore, we get that any eigenvector $v\in\mathop{\mathbb{R}}^{n}$ of $\mathcal{N}(M_{t})$, with eigenvalue $\lambda$, is also an eigenvector of $I-X_{t}^{4\delta}$ with eigenvalue at least $\frac{1}{3}\lambda$. The result follows since both matrices are symmetric and therefore have a spanning basis of eigenvectors. ∎
A.2 Fair Cuts
In this section, we prove Lemma 3.6 and Theorem 5.1, concerning the flow routines in the cut-matching step and the trimming step. The proofs are based on fair cuts, a stronger notion of approximate min-cut, introduced by [7].
Definition A.1 (Fair Cut, [7, 6]).
Let $G=(V,E)$ be an undirected graph with edge capacities $c\in\mathop{\mathbb{R}}^{E}_{>0}$. Let $s,t$ be two vertices in $V$. For any parameter $\alpha\geq 1$, we say that a cut $(S,T)$ is a $\alpha$-fair $(s,t)$-cut if there exists a feasible $(s,t)$-flow $f$ such that $f(u,v)\geq\frac{1}{\alpha}\cdot c(u,v)$ for every $(u,v)\in E(S,T)$, where $u\in S$ and $v\in T$.
Theorem A.2 (Fair Cut, Theorem 1.1 of [7]).
Given a graph $G=(V,E)$, two vertices $s,t\in V$ and $\epsilon\in(0,1]$, we can compute with high probability a $(1+\epsilon)$-fair $(s,t)$-cut, together with a corresponding feasible flow $f$, in $\tilde{O}(\frac{m}{\epsilon^{3}})$ time.
Proof of Lemma 3.6.
Let $\alpha=0.1$. Consider the following $(s,t)$-flow problem on a new graph $H=(V_{H}=V\cup\{s,t\},E_{H},w_{H})$ constructed as follows. We begin with $G$ and set its edge capacities to $\frac{c}{1+\alpha}$. We then add the source $s$ and connect it to every $v\in A^{l}$ with capacity $m_{v}$. Finally, we add the sink $t$ and connect it to every $v\in A^{r}$ with capacity $\frac{1}{1+\alpha}\bar{m}_{v}$.
We compute a $(1+\alpha)$-fair cut $(S,T)$ with a corresponding feasible flow $f$ in $\tilde{O}(m)$ time. We split into the following two cases:
Case 1: $S=\{s\}.$ By the definition of fair cut, for every $v\in A^{l}$, $f$ sends from $s$ to $v$ at least $\frac{1}{1+\alpha}m_{v}$ flow. By decomposing $f$ into flow paths and reducing the flow sent on each path, we obtain a flow $f^{\prime}$ for $H$ that sends from $s$ to $v$ exactly $\frac{1}{1+\alpha}m_{v}$ flow, for every $v\in A^{l}$. We claim that $f^{\prime\prime}=(1+\alpha)f^{\prime}$ is a feasible flow for $\Pi(G)$. Indeed, in $f^{\prime\prime}$, every $v\in A^{l}$ receives $m_{v}=$ flow from $s$, every $v\in A^{r}$ sends at most $(1+\alpha)\cdot\frac{1}{1+\alpha}\bar{m}_{v}=\bar{m}_{v}$ flow to $t$, and every $e\in E$ carries at most $(1+\alpha)\cdot\frac{1}{1+\alpha}c=c$ units of flow.
Case 2: $S\neq\{s\}.$ We split the cut edges $E_{H}(S,T)$ into three sets: (i) the edges of $G$ crossing the cut $E_{m}=E_{G}(S\setminus\{s\},T\setminus\{t\})$, (ii) the cut edges incident to $s$ $E_{s}=\{(s,v)\in E_{H}\mid v\in T\}$, and (iii) the cut edges incident to $t$ $E_{t}=\{(v,t)\in E_{H}\mid v\in S\}$. Let $\bar{E}_{s}$ be the incident edges to $s$ that are not in $E_{s}$.
We decompose $f$ into paths. By the definition of a fair cut, every edge $(s,v)\in E_{s}$ sends at least $\frac{1}{1+\alpha}m_{v}$ flow. We scale down the flow in the flow paths as follows. We delete all flow paths that start with an edge of $\bar{E}_{s}$. The rest of the flow paths are scaled down until each edge $(s,v)\in E_{s}$ sends exactly $\frac{1}{(1+\alpha)}m_{v}$ flow. This results in a flow $f^{\prime}$. The algorithm then returns $f^{\prime\prime}=(1+\alpha)f^{\prime}$ and the cut $S^{\prime}=S\setminus\{s\}$.
Similarly to the previous case, $f^{\prime\prime}$ is a feasible flow for the flow problem $\Pi(G\setminus S^{\prime})$. It remains to prove that $\Phi_{G}^{\mu}(S^{\prime},V\setminus S^{\prime})\mathrel{\mathop{:}}=\frac{|E_{m}|}{\min\{\mu(S^{\prime}),\mu(V\setminus S^{\prime})\}}\leq\frac{7}{c}$. We begin by upper bounding the numerator. For the remainder of the analysis, we consider again the flow $f$ and its path decomposition. By the definition of a fair cut, every flow path that traverses through $E_{m}$ must start with an edge in $\bar{E}_{s}$, and every edge in $E_{m}$ is at least $\frac{1}{1+\alpha}$ saturated. Therefore
$\frac{c}{1+\alpha}\cdot|E_{G}(S^{\prime},V\setminus S^{\prime})|=w_{H}(E_{m})\leq(1+\alpha)w_{H}(\bar{E}_{s})=(1+\alpha)m(S^{\prime})\leq(1+\alpha)\mu(S^{\prime}).$ (5)
To bound the denominator we prove that $\mu(V\setminus S^{\prime})=\Omega(\mu(S^{\prime}))$. Recall the assumption of the lemma $\bar{m}(A^{r})\geq\mu(V)/2$. We prove that $\bar{m}(A^{r}\cap S^{\prime})$ is small. Indeed, for every $v\in A^{r}\cap S^{\prime}$, $(v,t)\in E_{t}$ is crossing the cut $(S,T)$ and therefore is $\frac{1}{1+\alpha}$ saturated. Since the total flow crossing $(S,T)$ is bounded by the total source capacity (which is $m(A^{l})$), we get that
$\frac{1}{1+\alpha}\bar{m}(A^{r}\cap S^{\prime})=w_{H}(E_{t})\leq(1+\alpha)m(A^{l})\leq\frac{1+\alpha}{8}\mu(V),$
where the last inequality follows from the assumption of the lemma. Putting everything together
$\mu(V\setminus S^{\prime})\geq\bar{m}(V\setminus S^{\prime})=\bar{m}(A^{r})-\bar{m}(A^{r}\cap S^{\prime})\geq\frac{\mu(V)}{2}-\frac{(1+\alpha)^{2}}{8}\mu(V)\geq\frac{1}{3}\mu(V)\geq\frac{1}{3}\mu(S^{\prime}).$
Together with (5), we conclude that
$\Phi_{G}^{\mu}(S^{\prime},V\setminus S^{\prime})\mathrel{\mathop{:}}=\frac{|E_{G}(S^{\prime},V\setminus S^{\prime})|}{\min\{\mu(S^{\prime}),\mu(V\setminus S^{\prime})\}}\leq\frac{\frac{(1+\alpha)^{2}}{c}\mu(S^{\prime})}{\frac{1}{3}\mu(S^{\prime})}\leq\frac{7}{c}.$
As a direct corollary, we conclude that if Lemma 3.6 terminates in case $2$, then there exists a flow of constant congestion that routes one unit of flow from every edge in the cut $E(S,V\setminus S)$ into $S$, such that each vertex $v\in S$ receives at most $O(\phi\log n\cdot\mu(v))$ flow.
Corollary A.3.
Let $(S,V\setminus S)$ be the cut returned by Lemma 3.6, and assume $S\neq\emptyset$. Then there exists a flow of congestion $O(1)$ in $G[S]$ such that every $v\in S$ sends $\left(deg_{G}-deg_{G[S]}\right)(v)$ flow and receives at most $O(\phi\log n\cdot\mu(v))$ flow.
Recall that during the cut-matching game, the matching player produces cuts $S_{1},\ldots,S_{T}$ by applying Lemma 3.6. Since these cuts are disjoint, we get the following lemma.
Lemma A.4.
Let $(A,\bar{A})$ be the cut returned by case $2$ of Theorem 5.2. Then there exists a flow of congestion $O(1)$ in $G[\bar{A}]$ such that every $v\in\bar{A}$ sends $\left(deg_{G}-deg_{G[\bar{A}]}\right)(v)$ flow and receives at most $O(\phi\log n\cdot\mu(v))$ flow.
Proof.
Follows from Corollary A.3. ∎
Proof of Theorem 5.1.
Assume $|E[A,\bar{A})|>0$, as otherwise $A$ is a $(\phi,\mu)$-expander and we set $A^{\prime}=A$. Consider the following $(s,t)$-flow problem on a new graph $H=(V_{H}=A\cup\{s,t\},E_{H})$ constructed as follows. Start from $G$, and contract $V\setminus A$ into a single vertex, labeled as the source $s$ (ignoring self loops at $s$). Next, set the capacity of each edge ($G[A]$ and edges touching $s$) to be $\frac{3}{\phi}$. Finally, add a new sink vertex $t$ and connect it to each vertex $v\in A$ with an edge of capacity $\mu(v)$. Let $\alpha=0.1$, and compute a $(1+\alpha)$-fair cut $(S,T)$ of $H$ in time $\tilde{O}(m)$. Let $A^{\prime}=T\setminus\{t\}$. [^16] We later show that $A^{\prime}\neq\emptyset$. We now show that $A^{\prime}$ satisfies the requirements of the lemma.
First, suppose for contradiction that $G[A^{\prime}]$ is not a $\frac{\phi}{6}$-expander with respect to $\mu$. Then, there is a violating set $U\subseteq A^{\prime}$, satisfying $0<\mu(U)\leq\mu(A^{\prime}\setminus U)$ and
$|E(U,A^{\prime}\setminus U)|<\frac{\phi}{6}\mu(U)~.$
Since $A$ is a near $(\phi,\mu)$-expander with respect to $\mu$ then so is $A^{\prime}$ and therefore,
$|E(U,V\setminus U)|\geq\phi\mu(U)~.$
Taking the difference of the two inequalities above,
$|E(U,V\setminus A^{\prime})|=|E(U,V\setminus U)|-|E(U,A^{\prime}\setminus U)|\geq\frac{5\phi}{6}\mu(U)~.$
Since $(S,T)$ is a $(1+\alpha)$-fair cut, there is a feasible flow $f$ that saturates each edge of $E_{H}(S,T)$ up to a factor of $\frac{1}{1+\alpha}$. Each edge $(u,v)$ in $E(U,V\setminus A^{\prime})$ corresponds to an edge in $E_{H}(S,T)$ of capacity $\frac{3}{\phi}$, and the flow $f$ must send at least $\frac{1}{1+\alpha}\cdot\frac{3}{\phi}\geq\frac{2}{\phi}$ flow along the edge (in the direction from $S$ to $T$). In total, the amount of flow entering $U$ in $H$ is at least
$\frac{2}{\phi}|E(U,V\setminus A^{\prime})|\geq\frac{2}{\phi}\cdot\frac{5\phi}{6}\mu(U)=\frac{5}{3}\mu(U)~.$
On the other hand, at most $\mu(U)$ flow can leave $U$ along the edges incident to $t$, and at most
$\frac{3}{\phi}|E(U,A^{\prime}\setminus U)|\leq\frac{3}{\phi}\cdot\frac{\phi}{6}\mu(U)=\frac{1}{2}\mu(U)$
flow can cross from $U$ to $A^{\prime}\setminus U$. This totals at most $\frac{3}{2}\mu(U)$ flow that can exit $U$, which is strictly less than the $\geq\frac{5}{3}\mu(U)$ flow that enters $U$ (as $\mu(U)>0$), a contradiction. Thus $G[A^{\prime}]$ is a $\frac{\phi}{6}$-expander.
Finally, we show the properties $\mu(A^{\prime})\geq\mu(A)-\frac{4}{\phi}|E(A,\bar{A})|$ and $|E(A^{\prime},\bar{A^{\prime}})|\leq 2|E(A,\bar{A})|$ promised by the lemma. Since $(S,T)$ is a $(1+\alpha)$-fair cut, it is in particular a $(1+\alpha)$-approximate $(s,t)$-mincut. Since $(\{s\},V_{H}\setminus\{s\})$ is an $(s,t)$-cut of capacity $\frac{3}{\phi}|E(A,\bar{A})|$, it follows that the cut $(S,T)$ has capacity at most $(1+\alpha)\frac{3}{\phi}\cdot|E(A,\bar{A})|$. To establish the first inequality of the theorem, observe that every vertex $v\in A\setminus A^{\prime}$ lies on the $S$-side of the cut $(S,T)$, and therefore contributes $\mu(v)$ to the capacity of $(S,T)$ through the edge $(v,t)$. Summing over all $v\in A\setminus A^{\prime}$, we obtain
$\mu(A\setminus A^{\prime})\leq|E(S,T)|\leq(1+\alpha)\cdot\frac{3}{\phi}|E(A,\bar{A})|\leq\frac{4}{\phi}|E(A,\bar{A})|~,$
which proves the first property. Moreover, since $\mu(A)\geq\frac{9|E(A,\bar{A})|\cdot}{\phi}$, it follows that $A^{\prime}\neq\emptyset$. For the second inequality, note that each edges $(u,v)$ in $E(A^{\prime},\bar{A^{\prime}})$ corresponds to an edge in $E(S,T)$ with capacity $\frac{3}{\phi}$, so summing over all such edges
$\frac{3}{\phi}|E(A^{\prime},\bar{A^{\prime}})|\leq|E(S,T)|\leq(1+\alpha)\frac{3}{\phi}|E(A,\bar{A})|~,$
meaning $|E(A^{\prime},\bar{A^{\prime}})|\leq(1+\alpha)|E(A,\bar{A})|$ which proves the second property.
From the auxiliary flow problem defined in the preceding proof, we deduce the following corollary.
Corollary A.5.
Let $(A^{\prime},\bar{A^{\prime}})$ be the cut returned from the trimming step (Theorem 5.1). The following flows exist.
  1. 1.
    A flow of congestion $\frac{3}{\phi}$ in $G[A^{\prime}]$, such that every $v\in A^{\prime}$ sends $\frac{2}{\phi}\left(deg_{G}-deg_{G[A^{\prime}]}\right)(v)$ flow and receives at most $\mu(v)$ flow. That is, every edge in the cut $E(A^{\prime},\bar{A^{\prime}})$ sends $\frac{2}{\phi}$ flow and every vertex $v\in A^{\prime}$ receives at most $\mu(v)$ flow.
  2. 2.
    A flow of congestion $O(1)$ in $G[\bar{A^{\prime}}]$ such that every $v\in\bar{A}$ sends $\left(deg_{G}-deg_{G[\bar{A}]}\right)(v)$ flow and receives at most $O(\phi\log n\cdot\mu(v))$ flow. That is, every edge in the cut $E(A^{\prime},\bar{A^{\prime}})$ sends one unit of flow and every vertex $v\in\bar{A}$ receives at most $O(\phi\log n\mu(v))$ flow. [^17] Note that vertices in $\bar{A^{\prime}}\setminus\bar{A}$ don’t receive any flow.
Proof.
Let $f$ denote the flow constructed in the proof of Theorem 5.1. We decompose $f$ into flow paths from $s$ to $t$ and scale down each flow path so that the flow crossing every edge in the fair cut $E_{H}(S,T)$ equals $\frac{2}{\phi}$. Let $f^{\prime}$ denote the resulting flow.
For the first item, we delete from $f^{\prime}$ all flow paths that contain an edge $(v,t)$, where $t$ is the sink and $v\in S=\bar{A^{\prime}}\setminus\bar{A}$. Denote the resulting flow $f^{\prime\prime}$. The requested flow is achieved by restricting the flow paths in $f^{\prime\prime}$ to their subpath that intersects $T\setminus\{t\}$ (that is, deleting the prefix in $S$ and the last vertex $t$).
We proceed to the second item, and prove it in two steps. We first route flow from the new cut edges $E_{G}(A^{\prime},\bar{A^{\prime}})\setminus E_{G}(A,\bar{A})$ to $E_{G}(A,\bar{A})$, and then we route the flow from $E_{G}(A,\bar{A})$ to $\bar{A}$ so that every $v\in\bar{A}$ receives at most $O(\phi\log n\mu(v))$ flow.
For the first step, observe that there is a flow $f_{1}$ of congestion $1.5=O(1)$ in $G[\bar{A^{\prime}}]$ such that every $v\in\bar{A^{\prime}}\setminus\bar{A}=S\setminus\{s\}$ sends $(deg_{G}-deg_{G[\bar{A^{\prime}}]})(v)$ flow and receives at most $1.5(deg_{G}-deg_{G[\bar{A}]})(v)$ flow. This flow is obtained by scaling the flow paths of $f^{\prime}$ by $\phi/2$ and then restricting them to their prefix from $s$ until the first edge in $T$.
In the second step, note that the cut $(A,\bar{A})$ was produced in the cut-matching step. Therefore, by Corollary A.3, there exists a flow $f_{2}$ of congestion $O(1)$ in $G[\bar{A}]$ such that every $v\in\bar{A}$ sends $(deg_{G}-deg_{G[\bar{A}]})(v)$ flow and receives at most $O(\phi\log n\cdot\mu(v))$ flow.
The final flow is as follows: We first route according to $f_{1}$. Each $v\in\bar{A}$ must then send a total of $2.5(deg_{G}-deg_{G[\bar{A^{\prime}}]})(v)$ flow: this consists of $(deg_{G}-deg_{G[\bar{A^{\prime}}]})(v)$ flow required by the lemma and an additional $1.5(deg_{G}-deg_{G[\bar{A^{\prime}}]})(v)$ flow from $f_{1}$. We route this flow using a scaled version of $f_{2}$, which incurs at most $2.5\cdot O(1)$ congestion. ∎
Appendix B Algebraic Tools
Fact B.1 (Exercise IX.3.3 [4]).
Let $X,Y,A\in\mathop{\mathbb{R}}^{n\times n},m\in\mathop{\mathbb{N}}$, then
  1. 1.
    $\operatorname{Tr}(XY)=\operatorname{Tr}(YX)$.
  2. 2.
    $\operatorname{Tr}(A^{2m})\leq\operatorname{Tr}((A\cdot A^{\prime})^{m})$.
Proof.
  1. 1.
    Follows since $\operatorname{Tr}(XY)=\sum_{i=1}^{n}{\sum_{j=1}^{n}{X(i,j)Y(j,i)}}=\sum_{j=1}^{n}{\sum_{i=1}^{n}{Y(j,i)X(i,j)}}=\operatorname{Tr}(YX)$.
  2. 2.
    Using Schur decomposition, we decompose $A=VUV^{*}$, where $V\in\mathop{\mathbb{C}}^{n\times n}$ is a unitary matrix and $U\in\mathop{\mathbb{C}}^{n\times n}$ is an upper triangular matrix. Observe that
    $\operatorname{Tr}(A^{2m})=\operatorname{Tr}((VUV^{*})^{2m})=\operatorname{Tr}(VU^{2m}V^{*})=\operatorname{Tr}(V^{*}VU^{2m})=\operatorname{Tr}(U^{2m})=\sum_{i=1}^{n}{U(i,i)^{2m}}.$
    Let $S=UU^{*}$, then $S(i,i)=\sum_{j=1}^{n}{|U(i,j)|^{2}}\geq|U(i,i)|^{2}$, for every $i$. Moreover, $S$ is PSD Hermetian matrix so $S$ can be decomposed to $S=RDR^{*}$, where $R$ is a unitary matrix and $D$ is diagonal with non-negative real entries. Since $R$ is unitary, it holds that $\sum_{j=1}^{n}|R(i,j)|^{2}=1$, for every $i$. By convexity, we get
    $\left(S^{m}\right)(i,i)$ $=\left(RD^{m}R^{*}\right)(i,i)=\sum_{j=1}^{n}{|R(i,j)|^{2}D(j,j)^{m}}\geq\left(\sum_{j=1}^{n}{|R(i,j)|^{2}D(j,j)}\right)^{m}$
    $=(S(i,i))^{m}\geq|U(i,i)|^{2m}.$
    Therefore
    $\operatorname{Tr}((A\cdot A^{\prime})^{m})$ $=\operatorname{Tr}((VUU^{*}V^{*})^{m})=\operatorname{Tr}(V(UU^{*})^{m}V^{*})=\operatorname{Tr}((UU^{*})^{m})$
    $=\operatorname{Tr}(S^{m})=\sum_{i=1}^{n}\left(S^{m}\right)(i,i)\geq\sum_{i=1}^{n}(U(i,i))^{2m}=\operatorname{Tr}(A^{2m}).$
Lemma B.2 (In the proof of Theorem IX.3.5 [4]).
Let $X,Y\in\mathop{\mathbb{R}}^{n\times n}$ be symmetric matrices. Then for any positive integer $k$, $\operatorname{Tr}((XY)^{2^{k}})\leq\operatorname{Tr}(X^{2^{k}}Y^{2^{k}})$.
Proof.
By Fact B.1, we get
$\operatorname{Tr}((XY)^{2^{k}})\leq\operatorname{Tr}(((XY)(XY)^{\prime})^{2^{k-1}})=\operatorname{Tr}((XY^{2}X)^{2^{k-1}})=\operatorname{Tr}((X^{2}Y^{2})^{2^{k-1}}).$
The result follows by induction. ∎
Theorem B.3 (Symmetric Rearrangement; Theorem A.2 [9]).
Let $X,Y\in\mathop{\mathbb{R}}^{n\times n}$ be symmetric matrices. Then for any positive integer $k$,
$\operatorname{Tr}\left((XYX)^{2^{k}}\right)\leq\operatorname{Tr}\left(X^{2^{k}}Y^{2^{k}}X^{2^{k}}\right)\ .$
Proof.
By Lemma B.2 and Fact B.1
$\operatorname{Tr}\left((XYX)^{2^{k}}\right)=\operatorname{Tr}((X^{2}Y)^{2^{k}})\leq\operatorname{Tr}(X^{2^{k+1}}Y^{2^{k}})=\operatorname{Tr}(X^{2^{k}}Y^{2^{k}}X^{2^{k}})$
The following lemma is standard.
Lemma B.4.
For every weighted graph $G=(V,E,w)$ and $v\in\mathbb{R}^{n}$, where $n=|V|$, it holds that $v^{t}\mathcal{L}(G)v=\sum_{(s,t)\in E}{w(s,t)\cdot(v_{s}-v_{t})^{2}}$.
Lemma B.5.
Let $G=(V=[n],E,w)$ be a weighted undirected graph. For every $A\in\mathop{\mathbb{R}}^{n\times n}$, it holds that $\operatorname{Tr}(A^{\prime}\mathcal{L}(G)A)=\sum_{\{i,j\}\in E}{w(i,j)\cdot\left\|A(i)-A(j)\right\|_{2}^{2}}$, where $A(i)\in\mathop{\mathbb{R}}^{n}$ is row $i$ of $A$.
Proof.
The $i$’th column of $A$ is $A^{\prime}(i)\in\mathop{\mathbb{R}}^{n}$. By Lemma B.4
$\operatorname{Tr}(A^{\prime}\mathcal{L}(G)A)$ $=\sum_{k=1}^{n}{\left(A^{\prime}(k)\right)^{\prime}\mathcal{L}(G)A^{\prime}(k)}=\sum_{k=1}^{n}{\sum_{\{i,j\}\in E}{w(i,j)\cdot(A(i,k)-A(j,k))^{2}}}$
$=\sum_{\{i,j\}\in E}{\sum_{k=1}^{n}{w(i,j)\cdot(A(i,k)-A(j,k))^{2}}}=\sum_{\{i,j\}\in E}{w(i,j)\cdot\left\|A(i)-A(j)\right\|_{2}^{2}}$
Lemma B.6.
Let $A\in\mathop{\mathbb{R}}^{n\times n}$ be a symmetric $\mu$-stochastic $\mathcal{T}$-blocked matrix, then the eigenvalues of the normalized Laplacian $\mathcal{N}(A)=I_{\mathcal{T}}-U^{-\frac{1}{2}}AU^{-\frac{1}{2}}$ are in the range $[0,2]$.
Proof.
Since $\mathcal{L}(A)\geq 0$ we get that $\mathcal{N}(A)=U^{-\frac{1}{2}}\mathcal{L}(A)U^{-\frac{1}{2}}\geq 0$. Therefore, it is left to show that the eigenvalues of $U^{-\frac{1}{2}}AU^{-\frac{1}{2}}$ are at least 1-1. Let $X=U+A$. Note that $X\geq 0$. Indeed, similarly to Lemma B.4 ,for any $v\in\mathop{\mathbb{R}}^{n}$
$v^{\prime}Xv=\sum_{i=1}^{n}\sum_{j=1}^{n}A(i,j)(v_{i}+v_{j})^{2}\geq 0.$
Hence, $I_{\mathcal{T}}+U^{-\frac{1}{2}}AU^{-\frac{1}{2}}=U^{-\frac{1}{2}}XU^{-\frac{1}{2}}\geq 0$ which means that the eigenvalues of $U^{-\frac{1}{2}}AU^{-\frac{1}{2}}$ are at least 1-1. ∎
Lemma B.7.
Let $\{v_{i}\}_{i=1}^{k}$ be a set of $k$ vectors in $\mathop{\mathbb{R}}^{n}$. Then,
$k\left\|\frac{\sum_{i=1}^{k}{v_{i}}}{k}\right\|_{2}^{2}\leq\sum_{i=1}^{k}{\left\|v_{i}\right\|_{2}^{2}}$
Proof.
By the Cauchy-Schwartz inequality, for all $j\in[n]$,
$k\left({\frac{\sum_{i=1}^{k}{v_{i,j}}}{k}}\right)^{2}=\frac{1}{k}\left({\sum_{i=1}^{k}{v_{i,j}}}\right)^{2}\leq\frac{1}{k}\cdot k\cdot\sum_{i=1}^{k}{v_{i,j}^{2}}=\sum_{i=1}^{k}{v_{i,j}^{2}}$
Therefore,
$k\left\|\frac{\sum_{i=1}^{k}{v_{i}}}{k}\right\|_{2}^{2}=\sum_{j=1}^{n}{k\left({\frac{\sum_{i=1}^{k}{v_{i,j}}}{k}}\right)^{2}}\leq\sum_{j=1}^{n}{\sum_{i=1}^{k}{v_{i,j}^{2}}}=\sum_{i=1}^{k}{\sum_{j=1}^{n}{v_{i,j}^{2}}}=\sum_{i=1}^{k}{\left\|v_{i}\right\|_{2}^{2}}$
Appendix C Projection Lemmas
The following fact is quite standard in the analysis of algorithms in the cut-matching framework.
Lemma C.1 (Gaussian Behavior of Projections; Variation of Lemma 3.5 [5]).
Let $\{v_{i}\}_{i=1}^{k}$ be a set of $k\leq n$ vectors in $\mathop{\mathbb{R}}^{n}$. For $i\in[k]$, let $u_{i}=\langle v_{i},r\rangle$ be the projection of $v_{i}$ onto a random unit vector $r\in\mathbb{S}^{n-1}\subseteq\mathop{\mathbb{R}}^{n}$. Then:
  1. 1.
    $\mathop{\mathbb{E}}[u_{i}^{2}]=\frac{1}{n}\left\|v_{i}\right\|_{2}^{2}$ for all $i$.
  2. 2.
    $u_{i}^{2}\leq\frac{\alpha\log n}{n}\left\|v_{i}\right\|_{2}^{2}$
    holds for all $i$ with probability of at least $1-\frac{1}{n^{\alpha/8}}$, for any $\alpha\geq 8$.
Remark C.2.
Lemma C.1 follows from Lemma 3.5 of [5] as follows: For a single projection, condition (2) fails to hold with probability at most $e^{-\alpha\log n/4}=n^{-\alpha/4}$. By the union bound it fails for some vector with probability $n^{1-\alpha/4}$. For this failure probability to be small we need $\alpha$ to be sufficiently large; If $\alpha\geq 8$ then the failure probability is at most $n^{-\alpha/8}$.
Corollary C.3.
Let $\{v_{i}\}_{i=1}^{k}$ be a set of $k\leq n$ vectors in $\mathop{\mathbb{R}}^{n}$. For $i\in[k]$, let $u_{i}=\langle v_{i},r\rangle$ be the projection of $v_{i}$ onto a random unit vector $r\in\mathbb{S}^{n-1}\subseteq\mathop{\mathbb{R}}^{n}$. Then:
  1. 1.
    $\mathop{\mathbb{E}}[u_{i}^{2}]=\frac{1}{n}\left\|v_{i}\right\|_{2}^{2}$ for all $i$, and $\mathop{\mathbb{E}}[(u_{i}-u_{j})^{2}]=\frac{1}{n}\left\|v_{i}-v_{j}\right\|_{2}^{2}$ for all pairs $(i,j)$.
  2. 2.
    $u_{i}^{2}$ $\leq\frac{\alpha\log n}{n}\left\|v_{i}\right\|_{2}^{2}$
    $(u_{i}-u_{j})^{2}$ $\leq\frac{\alpha\log n}{n}\left\|v_{i}-v_{j}\right\|_{2}^{2}$
    holds for all indices $i$ and pairs $(i,j)$ with probability of at least $1-\frac{1}{n^{\alpha/8}}$, for some constant $\alpha\geq 16$.
Proof.
Apply Lemma C.1 and Remark C.2 on the set of vectors
$\{v_{i}:i\in[k]\}\cup\{v_{i}-v_{j}:i,j\in[k]\}$